### v1

# Achronal localization and representation of the causal logic from conserved current, application to massive scalar boson

Abstract
Covariant achronal localizations are gained out of covariant conserved currents computing their flux passing through achronal surfaces. This general method applies to the probability density currents with causal kernel regarding the massive scalar boson. Due to the one-to-one correspondence between (covariant) achronal localizations and (covariant) representations of the causal logic thus, apparently for the first time, a covariant representation of the causal logic for an elementary relativistic quantum mechanical system has been achieved. Similarly one derives the covariant family of representations of the causal logic related to the stress energy tensor of the massive scalar boson.
While reaching this result the divergence theorem is proven for open sets with almost Lipschitz boundary.
1 Introduction
Achronal localization. There are plain physical grounds why in order to comply with causality localization of a relativistic quantum mechanical system has to occur in all achronal regions of spacetime. This means that to every achronal Borel set $\Delta\subset\mathbb{R}^{4}$ there is assigned a nonnegative bounded operator $T(\Delta)$ acting on the Hilbert space of states with the expectation value $\langle\phi,T(\Delta)\phi\rangle$ being the probability of localization in $\Delta$ of the system in the state $\phi$. As argued further in [6] the probabilities of localization regarding achronally separated regions add up and yield $1$ for every maximal achronal region. In summary $T$ is a positive operator valued normalized measure (POVM) on every maximal achronal set. Moreover, by relativistic symmetry, $T$ is Poincaré covariant with respect to the unitary kinematical transformations of the states.
       Localization on spacelike Cauchy surfaces. Just in order to meet the requirements of causal localization [11] extended localization from flat spacelike regions, as up to then commonly considered, to proper spacetime regions studying thoroughly POVM on spacelike smooth Cauchy surfaces. The POVM are coherent in the sense that the localization operators assigned to a region contained in the intersection of two different spacelike smooth Cauchy surfaces coincide. An important result is that the localization considered in [11] satisfies a rather general causality requirement. The considerations actually extend to $C^{1}$ Cauchy surfaces. Here we do the final step to achronal localization described above.
       The construction of an achronal localization goes along the lines of the construction of the Cauchy localization in [11]. However there are to overcome some technical difficulties due to the fact that the Lipschitz continuous function, with its graph being the achronal set under consideration, in general is not $C^{1}$.
       Divergence theorem. For this purpose we prove the divergence theorem for open bounded subsets of $\mathbb{R}^{n}$ with almost Lipschitz boundary such that the boundary has finite $(n-1)$-dimensional Hausdorff measure and the irregular points of the boundary are contained in a compact set of zero $(n-1)$-dimensional Minkowski content. This extension of the divergence theorem (3) is new. In view of applications we add some remarks about how verify the assumptions of the theorem.
       Flux through maximal achronal sets. Using the prior result the crucial result is proven that the future-directed flux passing through a spacelike Cauchy surface of a conserved bounded zero or causal future-directed $C^{1}$-current is the same for all these surfaces (4).
       In corollary (4) the result of (4) is extended to all maximal achronal sets. (4) is proven under an additional boundedness assumption. Regarding the application to the massive scalar boson this assumption turns out to be almost irrelevant.
       Construction of achronal localization. At this juncture one is ready to derive the main result (5). Roughly speaking, every covariant conserved $C^{1}$ current with bounded zeroth component being positive quadratic on Euclidean space determines by the flux through the achronal sets a covariant achronal localization.
       Application to the massive scalar boson. The localizations of the massive scalar boson in Euclidean space (i.e., the Euclidean covariant positive operator valued normalized measures in $\mathbb{R}^{3}$, called POL in [6]) are determined by the integrals of a positive definite rotational invariant probability density $J_{0}$ over the regions of localization [6, sec. 6]. $J_{0}$ is the zeroth component of a covariant conserved current $\mathfrak{J}$ if and and only if the kernel is causal (6). This result is by [13]. Causal kernels have been studied in [13], [14], [6]. Under the almost irrelevant condition that the causal kernel is $C^{4}$ it is easy to verify that $\mathfrak{J}$ satisfies the assumption in (5) thus giving rise to a covariant achronal localization of the massive scalar boson.
       Analogously one obtains a covariant family of covariant achronal localizations related to the stress energy tensor of the massive scalar boson [17].
       Representation of the causal logic. Every (covariant) achronal localization determines uniquely a (covariant) representation of the causal logic and vice versa [6, (21)]. Obviously this one-to-one correspondence emphasizes further the relevance of achronal localization.
       Hence for the first time a covariant representation of the causal logic for an elementary relativistic quantum mechanical system is achieved. Also a covariant family of representation of the causal logic is derived corresponding to the above mentioned family of achronal localizations related to the stress energy tensor.
2 Notations and notions
Vectors in $\mathbb{R}^{4}$ are denoted by $\mathfrak{x}=(x_{0},x)$ with $x:=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$. Let $\varpi:\mathbb{R}^{4}\to\mathbb{R}^{3}$ denote the projection $\varpi(\mathfrak{x}):=x$. Representing Minkowski spacetime by $\mathbb{R}^{4}$ the Minkowski product of $\mathfrak{a},\mathfrak{a}^{\prime}\in\mathbb{R}^{4}$ is given by $\mathfrak{a}\cdot\mathfrak{a}^{\prime}:=a_{0}a_{0}^{\prime}-aa^{\prime}$, where for vectors $a,a^{\prime}$ in $\mathbb{R}^{3}$ the scalar product $a_{1}a^{\prime}_{1}+a_{2}a^{\prime}_{2}+a_{3}a^{\prime}_{3}$ is denoted by $aa^{\prime}$. Often we use the notation $\mathfrak{a}^{\cdot 2}:=\mathfrak{a}\cdot\mathfrak{a}$.
       $\tilde{\mathcal{P}}=ISL(2,\mathbb{C})$ is the universal covering group of the Poincaré group. It acts on $\mathbb{R}^{4}$ as
$g\cdot\mathfrak{x}:=\mathfrak{a}+\varLambda(A)\mathfrak{x}\quad\text{ for }g=(\mathfrak{a},A)\in\tilde{\mathcal{P}},\,\mathfrak{x}\in\mathbb{R}^{4}$ (2.1)
where $\varLambda:SL(2,\mathbb{C})\to O(1,3)_{0}$ is the universal covering homomorphism onto the proper orthochronous Lorentz group. For short one writes $A\equiv(0,A)$, $\mathfrak{a}\equiv(\mathfrak{a},I_{2})$, and $A\cdot\mathfrak{x}=\varLambda(A)\mathfrak{x}$. For $M\subset\mathbb{R}^{4}$ and $g\in\tilde{\mathcal{P}}$ define $g\cdot M:=\{g\cdot\mathfrak{x}:\mathfrak{x}\in M\}$.
       The group operation on $\tilde{\mathcal{P}}$ reads $(\mathfrak{a},A)(\mathfrak{a}^{\prime},A^{\prime})=(\mathfrak{a}+A\cdot\mathfrak{a}^{\prime},AA^{\prime})$ with identity element $(0,I_{2})$ and inverse $(\mathfrak{a},A)^{-1}=(-A^{-1}\cdot\mathfrak{a},A^{-1})$.
       The fourvector $\mathfrak{z}\in\mathbb{R}^{4}\setminus\{0\}$ is called timelike, lightlike, causal if $|z_{0}|>|z|$, $|z_{0}|=|z|$, $|z_{0}|\geq|z|$, respectively. It is future-directed if $z_{0}>0$.
       The set $\mathfrak{a}+\mathbb{R}\mathfrak{z}$ for $\mathfrak{a},\mathfrak{z}\in\mathbb{R}^{4}$, $\mathfrak{z}\neq 0$ is called a line. The line is timelike, lightlike, causal, if so is $\mathfrak{z}$.
       A set $\Delta\subset\mathbb{R}^{4}$ is said to be spacelike if $|x_{0}-y_{0}|<|x-y|$ for $\mathfrak{x},\mathfrak{y}\in\Delta$, $\mathfrak{x}\neq\mathfrak{y}$.
       A set $\Delta\subset\mathbb{R}^{4}$ is said to be achronal if $|x_{0}-y_{0}|\leq|x-y|$ for $\mathfrak{x},\mathfrak{y}\in\Delta$. By definition $\Delta$ is maximal achronal if $\Delta$ in not properly contained in an achronal set. An achronal set is maximal achronal if and only if it meets every timelike line. Note also that a maximal achronal set is always closed.
3 Divergence theorem on open sets with almost Lipschitz boundary
The following is inspired by Maggi [16, Remark 9.5, Theorem 9.6]. As to the notation, for $x\in\mathbb{R}^{n}$, $z\in\mathbb{R}^{n-1}$, $s>0$ put $x^{\prime}:=(x_{1},\dots,x_{n-1})$ and $x=(x^{\prime},x_{n})$, and let $B_{s}(x):=\{y\in\mathbb{R}^{n}:|y-x|<s\}$, $C(x,s):=\{y\in\mathbb{R}^{n}:|y^{\prime}-x^{\prime}|<s,|y_{n}-x_{n}|<s\}$, $D(z,s):=\{y\in\mathbb{R}^{n-1}:|y-z|<s\}$. Given an outer measure $\omega$ on $\mathbb{R}^{n}$ and $X\subset\mathbb{R}^{n}$, the trace (restriction) of $\omega$ on $X$ is the outer measure $\omega|X$ on $\mathbb{R}^{n}$ given by $\omega|X(F):=\omega(F\cap X)$. $\mathcal{H}^{n-1}$ denotes the $n-1$-dimensional Hausdorff measure on $\mathbb{R}^{n}$.
(1) Definition. Let $E\subset\mathbb{R}^{n}$ be open. $E$ has almost Lipschitz boundary $\partial E$ if there is a localization of $\partial E$ as follows. There is a closed set $M_{0}\subset\partial E$ with $\mathcal{H}^{n-1}(M_{0})=0$ such that for every $x\in M:=\partial E\setminus M_{0}$ there exists $s\in]0,\infty[$ with, up to rotation [^1] This means that there exists an orthonormal basis $e_{1},\dots,e_{n}$ of $\mathbb{R}^{n}$ such that (a),(b) hold for the coordinates of $x$ with respect to this basis.,
  • (a)
    $C(x,s)\cap E=\{y\in C(x,s):y_{n}>u(y^{\prime})\}$
  • (b)
    $C(x,s)\cap\partial E=C(x,s)\cap M=\{y\in C(x,s):y_{n}=u(y^{\prime})\}$
for some Lipschitz function $u:D(x^{\prime},s)\to\mathbb{R}$. Being $G$ the set of points of differentiability of $u$ define the unit vector field
$\nu:\operatorname{graph}u|_{G}\to\mathbb{R}^{n},\quad\nu(z,u(z)):=\frac{(\nabla u(z),-1)}{|(\nabla u(z),-1)|}$
Recall that by Rademacher’s theorem $G$ is the complement of a Lebesgue null set and that $\nabla u$ is measurable.
Henceforth $E\subset\mathbb{R}^{n}$ is open with almost Lipschitz boundary. The notation refers to (3).
(2) Lemma. $\mathcal{H}^{n-1}|C(x,s)\cap M$ is a Radon measure and $\mathcal{H}^{n-1}\big{(}(C(x,s)\cap M)\setminus\operatorname{graph}u|_{G}\big{)}=0$.
Proof.
Put $D:=D(x^{\prime},s)\cap\{z\in\mathbb{R}^{n-1}:|u(z)-x_{n}|<s\}$. Note $\operatorname{graph}\,u|_{D}=C(x,s)\cap\operatorname{graph}\,u=C(x,s)\cap\partial E=C(x,s)\cap M$.
       $\mathcal{H}^{n-1}|C(x,s)\cap M=\mathcal{H}^{n-1}|\operatorname{graph}\,u|_{D}$ is a Radon measure due to the area formula ([16, Theorem 9.1]) . Moreover, one has $\mathcal{H}^{n-1}(\operatorname{graph}\,u|_{D(x^{\prime},s)\setminus G})=\int_{D(x^{\prime},s)\setminus G}|(\nabla u(z),-1)|\operatorname{d}z=0$. ∎
The following preparatory result (3) comprises by the case $M_{0}=\emptyset$ the divergence theorem on bounded open sets with Lipschitz boundary (see also e.g. [16, Remark 9.5] and [1, A8.8]).
(3) Proposition. Let $E$ be bounded. Let $\varphi\in C_{c}^{1}(\mathbb{R}^{n})$ vanish on a neighborhood of $M_{0}$. Then
$\int_{E}\nabla\varphi\;\operatorname{d}\mathcal{L}^{n}=\int_{\partial E}\varphi\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}$
holds. Here $\nu_{E}$ is a $\mathcal{H}^{n-1}$-a.e. determined unit vector field on $\partial E$. Locally $\nu_{E}$ coincides with $\nu$ in (3).
Proof. (a) Assume first $\varphi\in C_{c}^{1}(C(x,s))$. By (3) the surface integral is well-defined. Following step one of the proof of [16, Theorem 9.3] the result follows due to $C(x,s)\cap\partial E=C(x,s)\cap M$ by definition (3)(b). Note that the weak gradient $\nabla f_{\delta}$ equals $\frac{1}{2\delta}(-\nabla u(z),1)$ at $z\in G$.
       (b) Now the vector field $\nu_{E}$ is constructed. Applying the result in (a) to all $\varphi\in C_{c}^{1}(C(x,s)\cap C(\bar{x},\bar{s}))$ shows that $\nu$ and $\bar{\nu}$ coincide a.e. on their common domain. There are countably many $C(x_{m},s_{m})$ from the localization in (3), which cover $M$. (Indeed, $M_{0}$ is closed and hence $G_{\delta}$. Since $\partial E$ is compact it follows that $M=\partial E\setminus M_{0}$ is $\sigma$-compact.) Thus we may compose from the corresponding $\nu_{m}$ by means of [16, Theorem 9.1] an $\mathcal{H}^{n-1}$-a.e. determined unit vector field $\nu_{E}$ on $\partial E$ which locally coincides a.e. with $\nu$ from (3).
       (c) As to the general case let $A\supset M_{0}$ be open with $\varphi|_{A}=0$. Note $\overline{E}\setminus A=(E\cup M)\setminus A\subset E\cup\bigcup_{x\in M}C(x,s)$. Since $\overline{E}\setminus A$ is compact it is covered by $E$ and finitely many $C(x,s)$. Hence one obtains a finite open cover of $\overline{E}$ by $U_{0}:=E$, $U_{k}$ being some $C(x,s)$ for $k=1,\dots,N$, and $U_{N+1}:=A$.
       Let $(\eta_{k})_{k=0,\dots,N+1}$ be a partition of unity for $\overline{E}$ subject to this cover [1, 4.20], i.e. $\eta_{k}\in C^{\infty}_{c}(\mathbb{R}^{n})$, $\operatorname{supp}\eta_{k}\subset U_{k}$, $\eta_{k}\geq 0$, and $\sum_{k=0}^{N+1}\eta_{k}(x)=1$ for $x\in\overline{E}$. Then
  • $\int_{E}\nabla(\eta_{0}\varphi)\operatorname{d}\mathcal{L}^{n}=0$ by the fundamental theorem of calculus, and obviously $\int_{\partial E}(\eta_{0}\varphi)\nu_{E}\operatorname{d}\mathcal{H}^{n-1}=0$
  • $1\leq k\leq N$: $\int_{E}\nabla(\eta_{k}\varphi)\operatorname{d}\mathcal{L}^{n}=\int_{\partial E}(\eta_{k}\varphi)\nu_{E}\operatorname{d}\mathcal{H}^{n-1}$ by (a), (b)
  • $k=N+1$: $\int_{E}\nabla(\eta_{N+1}\varphi)\operatorname{d}\mathcal{L}^{n}=0$ and $\int_{\partial E}(\eta_{N+1}\varphi)\nu_{E}\operatorname{d}\mathcal{H}^{n-1}=0$ as $\varphi|_{A}=0$
whence $\int_{E}\nabla\varphi\operatorname{d}\mathcal{L}^{n}=\sum_{k=1}^{N}\int_{E}\nabla(\eta_{k}\varphi)\operatorname{d}\mathcal{L}^{n}=\sum_{k=1}^{N}\int_{\partial E}(\eta_{k}\varphi)\nu_{E}\operatorname{d}\mathcal{H}^{n-1}=\int_{\partial E}\varphi\nu_{E}\operatorname{d}\mathcal{H}^{n-1}$.
       (d) In conclusion consider a further localization of $\partial E$ according (3) with $\tilde{\nu}_{E}$ the related unit vector field by (b). Then $\int_{\partial E}\varphi\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}=\int_{\partial E}\varphi\,\tilde{\nu}_{E}\;\operatorname{d}\mathcal{H}^{n-1}$ holds for all $\varphi\in C_{c}^{1}(\mathbb{R}^{n})$ with $\operatorname{supp}\varphi\cap M_{0}=\emptyset$, whence $\nu_{E}=\tilde{\nu}_{E}$ $\mathcal{H}^{n-1}$-a.e.∎
(4) Lemma. Let $X\subset\mathbb{R}^{n}$ be Borel. Then $\mathcal{H}^{n-1}|X$ is Borel regular.
Proof. Let $F\subset\mathbb{R}^{n}$. Since $\mathcal{H}^{n-1}$ is Borel regular, there are Borel sets $A,B\subset\mathbb{R}^{n}$ with $F\cap X\subset A$, $F\setminus X\subset B$ and $\mathcal{H}^{n-1}(F\cap X)=\mathcal{H}^{n-1}(A)$, $\mathcal{H}^{n-1}(F\setminus X)=\mathcal{H}^{n-1}(B)$. Then the Borel set $C:=(A\cap X)\cup(B\setminus X)$ satisfies $F\subset C$ with $F\cap X=A\cap X=C\cap X$ so that $\mathcal{H}^{n-1}(F\cap X)=\mathcal{H}^{n-1}(C\cap X)$, i.e., $\mathcal{H}^{n-1}|X(F)=\mathcal{H}^{n-1}|X(C)$. ∎
(5) Theorem. Let $E$ be open bounded with almost Lipschitz boundary. Assume $\mathcal{H}^{n-1}(\partial E)<\infty$ and the Minkowski content $\mathcal{M}^{n-1}(M_{0})=0$. Then $\mathcal{H}^{n-1}|\partial E$ is a Radon measure and for $\varphi\in C_{c}^{1}(\mathbb{R}^{n})$
$\int_{E}\nabla\varphi\;\operatorname{d}\mathcal{L}^{n}=\int_{\partial E}\varphi\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}$
holds. Here $E$ can be replaced by $\overline{E}$ as $\mathcal{L}^{n}(\partial E)=0$, and $\partial E$ can be replaced by $M$ as $\mathcal{H}^{n-1}(M_{0})=0$.
Proof.
$\mathcal{H}^{n-1}|\partial E$ is a Radon measure by (3). Let $\delta>0$ and let $A_{\delta}$ denote the $\delta$-neighborhood of $M_{0}$. Then
  • $\mathcal{H}^{n-1}(\partial E\cap A_{\delta})\to 0$ for $\delta\to 0$
  • $\delta^{-1}\mathcal{L}^{n}(A_{\delta})\to 0$ for $\delta\to 0$; in particular, $\mathcal{L}^{n}(A_{\delta})\to 0$
Indeed, the first claim holds as $\mathcal{H}^{n-1}(\partial E\cap A_{\delta})<\infty$ and $A_{\delta}\downarrow_{\delta}M_{0}$. The very definition of the Minkowski content [12, 3.2.27] implies the second claim.
       Furthermore by [1, 4.19] there is $f\in C^{\infty}_{c}(\mathbb{R}^{n})$ with $\operatorname{supp}f\subset A_{\delta}$, $0\leq f\leq 1$, $f|_{M_{0}}=1$, and $|\nabla f|\leq C/\delta$, where the finite constant $C$ does not depend on $\delta$.
       Note that (3) applies to $(1-f)\varphi$. So $\int_{E}\nabla\varphi\operatorname{d}\mathcal{L}^{n}=\int_{\partial E}(1-f)\varphi\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}+\int_{E}\nabla(f\varphi)\operatorname{d}\mathcal{L}^{n}=\int_{\partial E}\varphi\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}-\int_{\partial E}f\varphi\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}+\int_{E}(\nabla f)\varphi\operatorname{d}\mathcal{L}^{n}+\int_{E}f\nabla\varphi\operatorname{d}\mathcal{L}^{n}$. The last three summands vanish as $\delta\to 0$.
       Indeed, $|\int_{\partial E}f\varphi\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}|\leq\int_{\partial E}1_{A_{\delta}}|\varphi|\operatorname{d}\mathcal{H}^{n-1}\leq||\varphi||_{\infty}\mathcal{H}^{n-1}(\partial E\cap A_{\delta})\to 0$. Next, the crucial one, $|\int_{E}(\nabla f)\varphi\operatorname{d}\mathcal{L}^{n}|\leq\int_{E}C\delta^{-1}1_{A_{\delta}}|\varphi|\operatorname{d}\mathcal{L}^{n}=||\varphi||_{\infty}C\delta^{-1}\mathcal{L}^{n}(A_{\delta})\to 0$. Finally one has $|\int_{E}f\nabla\varphi\operatorname{d}\mathcal{L}^{n}|\leq\int_{E}1_{A_{\delta}}||\nabla\varphi||\operatorname{d}\mathcal{L}^{n}\leq||\nabla\varphi||_{\infty}\mathcal{L}^{n}(A_{\delta})\to 0$, thus accomplishing the proof.∎
An obviously equivalent formulation of (3) is
(6) Corollary. Let $E$ be as in (3) and $\mathfrak{v}\in C_{c}^{1}(\mathbb{R}^{n},\mathbb{R}^{n})$ be a vector field. Then
$\int_{E}\operatorname{div}\,\mathfrak{v}\;\operatorname{d}\mathcal{L}^{n}=\int_{\partial E}\mathfrak{v}\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}$
where $\operatorname{div}\,\mathfrak{v}:=\sum_{i=1}^{n}\partial_{i}v_{i}$ denotes the divergence of $\mathfrak{v}$ and $\mathfrak{v}\,\nu_{E}$ is the $\mathbb{R}^{n}$-scalar product of $\mathfrak{v}$ and $\nu_{E}$.
In view of an application of (3), (3) the following remarks on $\mathcal{H}^{n-1}(\partial E)<\infty$ and $\mathcal{M}^{n-1}(M_{0})=0$ may be useful. Note first that $\mathcal{M}^{n-1}(M_{0})=0$ implies $\mathcal{H}^{n-1}(M_{0})=0$.
       According to [16, Theorem 8.1], $\mathcal{H}^{n-1}(D)<\infty$ if $D$ is the image under an injective Lipschitz function on $\mathbb{R}^{n-1}$ in $\mathbb{R}^{n}$ of a Lebesgue measurable set $C$ of finite Lebesgue measure. If $C$ is a Lebesgue null set then $\mathcal{H}^{n-1}(D)=0$.
(7) Corollary. Let $\partial E$ be covered by finitely many sets $D_{i}$, each $D_{i}$ being the image under an injective Lipschitz function on $\mathbb{R}^{n-1}$ in $\mathbb{R}^{n}$ of a Lebesgue measurable $C_{i}$ of finite Lebesgue measure. Then $\mathcal{H}^{n-1}(\partial E)<\infty$.
If the images $D_{i}$ of countably many Lebesgue null sets $C_{i}$ cover $M_{0}$, then $\mathcal{H}^{n-1}(M_{0})=0$.
       According to [12, Theorem 3.2.39], $\mathcal{H}^{n-1}(D)=\mathcal{M}^{n-1}(D)=0$ holds, if $D\subset\mathbb{R}^{n}$ is closed and if there is a Lipschitz function $h:\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ mapping some bounded $C\subset\mathbb{R}^{n-1}$ onto $D$ or, equivalently, if $h$ is locally Lipschitz defined on $\overline{C}$ with $h(C)=D$. The latter holds because $\overline{C}$ is compact.
(8) Proposition. Let $M_{0}$ be the union of finitely many sets $D_{i}$ of Minkowski content $\mathcal{M}^{n-1}(D_{i})=0$. Then $\mathcal{M}^{n-1}(M_{0})=0$.
Moreover, $\mathcal{M}^{n-1}(D_{i})=0$ holds if $\mathcal{H}^{n-1}(D_{i})=0$ and if $D_{i}$ is the image of a compact set under a locally Lipschitz function.
Proof.
It remains to prove the first part of the assertion. The finite subadditivity of the upper Minkowski content yields $\mathcal{M}^{*n-1}(M_{0})\leq\sum_{i}\mathcal{M}^{*n-1}(D_{i})=\sum_{i}\mathcal{M}^{n-1}(D_{i})=0$ implying $\mathcal{M}^{n-1}(M_{0})=0$. ∎
4 Flux passing through a maximal achronal set
The result (4) essentially is an extension of [11, Proposition 37] to spacelike Cauchy surfaces. Note that [11, Proposition 37] is equally valid for a $C^{1}$ vector field in place of a smooth one. General spacetime theories study Cauchy surfaces. These are the sets which meet every inextendible timelike smooth curve exactly once [19, Chapter 14, Definition 28]. By [19, Chapter 14, Lemma 29] they meet even all inextendible causal smooth curves. According to [6, (9) Remark, (36) Theorem] a spacelike Cauchy surface is just a causal base, i.e., it is a spacelike [^2] A set is called spacelike according to sec. 2. set which intersects every causal line. It equals the graph of the corresponding $1$-Lipschitz function.
(9) Lemma. Let $\tau:\mathbb{R}^{3}\to\mathbb{R}$ and $S:=\{(\tau(x),x):x\in\mathbb{R}^{3}\}$. Then $S$ is a spacelike Cauchy surface if $|\tau(x)-\tau(y)|<|x-y|$ for $x\neq y$ and $\limsup_{|x|\to\infty}|\tau(x)|/|x|<1$ or if a fortiori $\tau$ is $L$-Lipschitz with $L<1$.
Proof. Obviously $S$ is spacelike. Assume $(\mathfrak{a}+\mathbb{R}\mathfrak{z})\cap S=\emptyset$ for some $\mathfrak{z}=(1,e)$ with $0<|e|\leq 1$. Let $a_{0}>\tau(a)$. (The case $a_{0}<\tau(a)$ is analogous.) Then by continuity $a_{0}+s>\tau(a+se)$ for all $s$. Hence for $s<-a_{0}$ one has $|\tau(a+se)|/|a+se|>|a_{0}+s|/|a+se|\to 1/|e|\geq 1$ for $s\to-\infty$. This contradicts the assumption $\limsup_{|x|\to\infty}|\tau(x)|/|x|<1$.∎
A $C^{1}$ vector field $\mathfrak{j}$ on $\mathbb{R}^{4}$ is said to satisfy the continuity equation if $\operatorname{div}\mathfrak{j}=0$, i.e., if
$\partial_{0}j_{0}(\mathfrak{x})+\partial_{1}j_{1}(\mathfrak{x})+\,\partial_{2}j_{2}(\mathfrak{x})+\partial_{3}j_{3}(\mathfrak{x})=0$ (4.1)
holds for $\mathfrak{x}\in\mathbb{R}^{4}$.
(10) Theorem. Let the $C^{1}$ vector field $\mathfrak{j}$ on $\mathbb{R}^{4}$ be bounded and satisfy the continuity equation (4.1). Suppose that it is zero or causal future-directed, i.e.
  • $j_{0}(\mathfrak{x})\geq|j(\mathfrak{x})|$ for all $\mathfrak{x}\in\mathbb{R}^{4}$
Then for every spacelike Cauchy surface $S$ being the graph of the corresponding $1$-Lipschitz function $\tau:\mathbb{R}^{3}\to\mathbb{R}$
$\int\big{(}j_{0}(\tau(x),x)-j(\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x=\int j_{0}(0,x)\operatorname{d}^{3}x$ (4.2)
holds. The integrands are nonnegative and the integrals may be infinite, $\nabla\tau$ is measurable a.e. determined.
Proof. By the assumption on $\mathfrak{j}$ and $|\nabla\tau|\leq 1$ the integrands are nonnegative. Let $k:\mathbb{R}^{3}\to\mathbb{R}$ be positive bounded integrable $C^{1}$ like $k(x)=(1+x^{2})^{-2}$.
       (a) The $C^{1}$ vector field $\mathfrak{v}=(v_{0},v):=(j_{0}+k,j)$ is bounded so that its flow is complete, satisfies the continuity equation, and everywhere holds $v_{0}>|v|$. Hence the integral curves $\gamma_{x}$ of $\mathfrak{v}$ determined by $\dot{\gamma}_{x}(s)=\mathfrak{v}(\gamma_{x}(s))$, $\gamma_{x}(0)=(0,x)$, $x\in\mathbb{R}^{3}$ are timelike future directed inextendible. So they meet $S$ just once, i.e., there is exactly one $\sigma(x)\in\mathbb{R}$ such that $\gamma_{x}(\sigma(x))\in S$. Vice versa the integral curve which starts at the point $\mathfrak{y}\in S$ meets $\{0\}\times\mathbb{R}^{3}$. This proves that $\mathfrak{h}:\mathbb{R}^{3}\to\mathbb{R}^{4}$, $\mathfrak{h}(x):=\gamma_{x}(\sigma(x))$ is injective with $\mathfrak{h}(\mathbb{R}^{3})=S$.
       (b) The claim is that $\mathfrak{h}$ is locally Lipschitz. By [11, Proposition 37], $\Phi:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}^{4}$, $\Phi(s,x):=\gamma_{x}(s)$ is a diffeomorphism. Write $\Phi=:(\varphi,\phi)$.
       From the implicit function theorem [15] applied to the function $f:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}$, $f:=\varphi-\tau\circ\phi$, it follows that $\sigma:\mathbb{R}^{3}\to\mathbb{R}$ is continuous and hence that $\mathfrak{h}$ is continuous. Indeed, since $f(s,x)=0$ is uniquely solved by $s=\sigma(x)$ for every $x\in\mathbb{R}^{3}$ it remains to verify that $f(\cdot,x)$ is injective for every $x\in\mathbb{R}^{3}$. Let $f(s,x)=f(s^{\prime},x)$. This implies $|\varphi(s,x)-\varphi(s^{\prime},x)|\leq|\phi(s,x)-\phi(s^{\prime},x)|$. Since $\gamma_{x}$ is timelike, $s=s^{\prime}$ follows.
       Now we show that $\mathfrak{h}$ is even locally Lipschitz. Put $\mathfrak{G}:=(\varphi,\operatorname{id}_{\mathbb{R}^{3}})$. As $\partial_{s}\varphi(s,x)=v_{0}(\Phi(s,x))>0$ one infers that $\mathfrak{G}$ is an diffeomorphisms. Check $g(h_{0}(x),x)=h(x)$ for $g:=\phi\circ\mathfrak{G}^{-1}$.
       Fix $x,y\in\mathbb{R}^{3}$. Define $\beta:[0,1]\to\mathbb{R}^{3}$, $\beta(u):=g(\mathfrak{z}(u))$ with $\mathfrak{z}(u):=\big{(}(h_{0}(y),y)-(h_{0}(x),x)\big{)}u+(h_{0}(x),x)$. Check $\beta(1)=h(y)$, $\beta(0)=h(x)$ and remind $\mathfrak{h}(\mathbb{R}^{3})=S$.
       Then $|h_{0}(y)-h_{0}(x)|=|\tau(h(y))-\tau(h(x))|\leq|h(y)-h(x)|$, whence it suffices to show that $h$ is locally Lipschitz, and further $|h(y)-h(x)|=|\beta(1)-\beta(0)|\leq|\beta^{\prime}(u_{0})|$ for some $u_{0}$ by the mean value theorem.
       Now $\beta^{\prime}(u)=D_{time}g(\mathfrak{z}(u))\big{(}(h_{0}(y)-h_{0}(x)\big{)}+D_{space}g(\mathfrak{z}(u))\big{(}y-x\big{)}$, whence $|h(y)-h(x)|\leq||D_{time}g(\mathfrak{z}(u_{0}))||\,|h(y)-h(x)|+|D_{space}g(\mathfrak{z}(u_{0}))|(y-x)$.
       Obviously $||D_{space}g||$ is locally bounded on $\mathbb{R}^{4}$. The time derivative yields $D_{time}g(\mathfrak{z})=(\frac{1}{v_{0}}v)(\mathfrak{G}^{-1}(\mathfrak{z}))$, whence $|D_{time}g(\mathfrak{z})|<1$ for all $\mathfrak{z}\in\mathbb{R}^{4}$. Hence $||D_{space}g||/\big{(}1-|D_{time}g|\big{)}$ is locally bounded on $\mathbb{R}^{4}$; for bounded $D\subset\mathbb{R}^{4}$ let $C_{D}<\infty$ be a bound on $D$.
       Let $K\subset\mathbb{R}^{3}$ be a compact ball. The image of $K$ under the continuous map $(h_{0},\operatorname{id}_{\mathbb{R}^{3}})$ is contained in a compact ball $D\subset\mathbb{R}^{4}$. In conclusion, $|h_{0}(y)-h_{0}(x)|\leq C_{D}|y-x|$ for all $x,y\in K$.
       (c) Let $0<R<\infty$. Put $A_{R}:=\{x\in\mathbb{R}^{3}:|x|=R\}$ and $B_{R}:=\{x\in\mathbb{R}^{3}:|x|<R\}$. Suppose that $\sigma(x)>0$ for $|x|<R$. Then $E_{R}:=\bigcup_{x\in B_{R}}\{\gamma_{x}(s):0<s<\sigma(x)\}\subset\mathbb{R}^{4}$ is open with $\partial E_{R}=M\cup M_{0}$, where $M:=\{0\}\times B_{R}\cup L_{R}\cup\mathfrak{h}(B_{R})$ with $L_{R}:=\bigcup_{x\in A_{R}}\{\gamma_{x}(s):0<s<\sigma(x)\}$, and $M_{0}:=\{0\}\times A_{R}\cup\mathfrak{h}(A_{R})$.
       One verifies that (3) applies to $n=4$, $E=E_{R}$ showing $\mathcal{H}^{3}(\partial E_{R})<\infty$ by (d), (e), and $\mathcal{M}^{3}(M_{0})=0$ by (f), (3).
       (d) Let $C\subset\mathbb{R}^{3}$ be bounded Borel. Then $\mathcal{H}^{3}(\{0\}\times C)<\infty$, $\mathcal{H}^{3}(\operatorname{graph}\tau|_{C})<\infty$ by [16, Theorem 8.1].
       (e) Let $D$ be a bounded Borel subset of the tube $T_{R}:=\{\gamma_{x}(s):s\in\mathbb{R},|x|=R\}$. Then $\mathcal{H}^{n-1}(D)<\infty$. Indeed, $A_{R}$ is covered by $A:=\{x\in A_{R}:|x_{1}|\leq R/3,|x_{2}|\leq R/3\}$ and finitely many rotations of $A$ around the origin. Hence it suffices to show that $\mathcal{H}^{n-1}(D^{\prime})<\infty$ for $D^{\prime}:=\{\gamma_{x}(s):|s|\leq S,x\in A\}$ and $0<S<\infty$.
       Recall the diffeomorphism $\Phi:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}^{4}$. Hence $f:=\mathbb{R}\times]-R/2,R/2[\times]-R/2,R/2[\to\mathbb{R}^{4}$, $f(s,x_{1},x_{2}):=\Phi\big{(}s;x_{1},x_{2},\sqrt{R^{2}-x_{1}^{2}-x_{2}^{2}}\,\big{)}$ is injective, $C^{1}$ whence Lipschitz on compact sets, and $f([-S,S]\times[-R/3,R/3]\times[-R/3,R/3])=D^{\prime}$. The claim holds by [16, Theorem 8.1].
       (f) Let $C\subset\mathbb{R}^{3}$ be a compact Lebesgue null set. Then $\mathcal{H}^{3}(\{0\}\times C)=0$ and $\mathcal{H}^{3}(\mathfrak{h}(C))=0$ hold by [16, Theorem 8.1]. Indeed, the former is obvious, the latter holds true as $\mathfrak{h}$ is Lipschitz on $C$ being compact.
       (g) Now let $\eta\in C_{c}^{1}(\mathbb{R}^{4})$ with $0\leq\eta\leq 1$ and $\eta|_{\overline{E}_{R}}=1$. Put $\phi:=\eta\mathfrak{v}$. Then (3) yields $0=\int_{M}\mathfrak{v}\,\nu_{E_{R}}\operatorname{d}\mathcal{H}^{3}$ as $\mathfrak{v}$ satisfies the continuity equation. Moreover the integration over $L_{R}$ yields 00 as the integrand is 00. So using [16, Theorem 9.1] it follows $\int_{\varpi(\mathfrak{h}(B_{R}))}\big{(}v_{0}(\tau(x),x)-v(\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}x^{3}=\int_{B_{R}}v_{0}(0,x)\operatorname{d}x^{3}$.
       (h) Finally, as explained in [11, Proposition 37], the condition $\sigma(x)>0$ for $|x|<R$ in (c) can be removed and the limit $R\to\infty$ carried out. One ends up with (4.2).∎
Under the assumptions made on the vector field $\mathfrak{j}$, the result in (4) states that the future-directed flux passing through a spacelike Cauchy surface is the same for all these surfaces. Under the additional assumption (4)(b) on $\mathfrak{j}$ the result in (4) is extended to all maximal achronal sets.
(11) Corollary. Let the $C^{1}$ vector field $\mathfrak{j}$ be bounded and satisfy the continuity equation (4.1). Suppose that for all $\mathfrak{x}\in\mathbb{R}^{4}$
  • (a)
    $j_{0}(\mathfrak{x})\geq|j(\mathfrak{x})|$,
  • (b)
    $\gamma\,j_{0}(\mathfrak{x})+k(x)\geq\,|j(\mathfrak{x})|$ for some nonnegative integrable bounded $C^{1}$ function $k:\mathbb{R}^{3}\to\mathbb{R}$ and some $10<\gamma<$ 1.
Then for every maximal achronal set $\Lambda$ being the graph of the corresponding $1$-Lipschitz function $\tau:\mathbb{R}^{3}\to\mathbb{R}$
$\int\big{(}j_{0}(\tau(x),x)-j(\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x=\int j_{0}(0,x)\operatorname{d}^{3}x$ (4.3)
holds. The integrands are nonnegative and the integrals may be infinite.
Proof. Due to (a) and $|\nabla\tau|\leq 1$ the integrands are nonnegative. Define the vector field $\mathfrak{l}=(l_{0},l)$ by $l_{0}(\mathfrak{x}):=\gamma j_{0}(\frac{1}{\gamma}x_{0},x)+k(x)$, $l(\mathfrak{x}):=j(\frac{1}{\gamma}x_{0},x)$.
       One easily checks $(\partial_{i}l_{i})(\mathfrak{x})=(\partial_{i}j_{i})(\frac{1}{\gamma}x_{0},x)$ for $i=0,1,2,3$. Hence $\mathfrak{l}$ satisfies the continuity equation.
       Obviously the function $\rho:=\gamma\tau$ is $\gamma$-Lipschitz. As $\gamma<1$, by (4) the graph of $\rho$ is a spacelike Cauchy surface.
       So (4) applies to $\mathfrak{l}$ and $\rho$, whence
$\int\big{(}l_{0}(\rho(x),x)-l(\rho(x),x)\nabla\rho(x)\big{)}\operatorname{d}^{3}x=\int l_{0}(0,x)\operatorname{d}^{3}x\>.$
Replace $l_{0},l$ and $\rho$ by their definitions and multiply both sides by $\frac{1}{\gamma}$. One obtains (4.3) up to the finite positive summand $\frac{1}{\gamma}\int k(x)\operatorname{d}^{3}x$ on both sides.∎
Obviously one has
(12) Lemma. The assumption (4)(b) is satisfied, if (4)(a) and $j_{0}(x_{0},x)\leq\frac{1}{1-\gamma}k(x)$ hold.
5 Covariant achronal localization out of covariant conserved current
Let $\mathcal{H}$ be a separable Hilbert space. Let $\mathcal{B}^{ach}$ denote the family of Borel subsets $\Delta$ of $\mathbb{R}^{4}$, which are achronal.
(13) Definition. Let $T(\Delta)$ for $\Delta\in\mathcal{B}^{ach}$ be a nonnegative bounded operator on $\mathcal{H}$. Suppose $T(\emptyset)=0$ and $\sum_{n}T(\Delta_{n})=I$ for every sequence $(\Delta_{n})$ of mutually disjoint sets in $\mathcal{B}^{ach}$ such that $\bigcup_{n}\Delta_{n}$ is maximal achronal. Then the map $T$ is called an achronal localization (AL).
       Let $W$ be a unitary representation of $\tilde{\mathcal{P}}$. Then the AL $T$ is said to be (Poincaré) covariant by means of $W$ if $T(g\cdot\Delta)=W(g)T(\Delta)W(g)^{-1}$ holds for $g\in\tilde{\mathcal{P}}$ and $\Delta\in\mathcal{B}^{ach}$.
       As mentioned the meaning of $T$ is that $\langle\phi,T(\Delta)\phi\rangle$ is the probability of localization of the quantum mechanical system in the spacetime region $\Delta$ if the system is in the state $\phi$.
       There exist AL with the localization operators $T(\Delta)$ being orthogonal projections [6, (22) Theorem]. In this case the localization operators commute. However, a quantum mechanical system localized by a projection valued AL necessarily does not have a semi-bounded energy operator. This no-go result following from Hegerfeldt ’s well-known theorem regards a first type of Einstein causality requirement whose modern generalized reformulation [4, 6] we shall present below.
       The notion of localization of the above definition seems appropriate to describe measurement processes where a quantum system is absorbed by the apparatus and no further localization measurments can performed on it. This is because, in case of subsequent measurements a second type of Einstein’s causality requirement essentially regarding the no-signaling condition has to be considered. To comply with this requirement, under some popular assumptions about the post measurement state, $T(\Delta)$ and $T(\Delta^{\prime})$ should commute if $\Delta$ and $\Delta^{\prime}$ cannot be joined by causal curves. This requirement cannot be fulfilled as a consequence of Malament’s theorem and its modern re-formulations [^3] A quick review on these issues related to various facets of Einstein’s causality and localization appears in the introduction of [17]..
       The first-type of causality requirement in the modern generalized perspective states that the probability of localization in a region of influence determined by the limiting velocity of light is not less than that in the region of actual localization. The spacetime region $\Delta^{\prime}\subset\mathbb{R}^{4}$ is a region of influence of $\Delta$ if all causal lines, which intersect $\Delta$, meet $\Delta^{\prime}$. Hence the condition imposed by causality on an AL reads
$T(\Delta)\leq T(\Delta^{\prime})\>.$ (CC)
In [11], (CC) condition was proved true in a special case where the sets $\Delta$ belong to smooth Cauchy surfaces and in [6], the final result has been established that an AL satisfies (CC) in full. As argued in [6], (CC) even necessitate achronal localization. In fact spacelike localization is not sufficient since (CC) induces the localization in achronal hyperplanes. This fact is reported in [6] and studied in detail in [7]. Hence we consider the study and the explicit construction of the achronal localization of the massive scalar boson to be of utmost relevance.
       It is the very principle of causality which let one think of the probability of localization as a conserved quantity reigned by an associated density current $\mathfrak{J}$. Indeed, by (5) a covariant AL can be constructed by means of a covariant conserved current assuming that on the Euclidean space its zeroth component describes the density of the probability of localization, namely explicitly $J_{0}(\phi;0,x)\geq 0$, $x\in\mathbb{R}^{3}$, and
$\langle\phi,T(\Delta)\phi\rangle=\int_{\Delta}J_{0}(\phi;0,x)\operatorname{d}^{3}x$ (5.1)
for Borel $\Delta\subset\mathbb{R}^{3}$.
(14) Definition. Let $\mathcal{D}$ be a $W$-invariant dense subspace of $\mathcal{H}$. Let $\mathfrak{J}=(J_{0},J)$ be a map from $\mathcal{D}\times\mathbb{R}^{4}$ to $\mathbb{R}^{4}$ such that $\mathfrak{J}(\phi,\cdot)$ is a bounded $C^{1}$ vector field. (i) $\mathfrak{J}$ is conserved if the latter satisfies the continuity equation. (ii) $\mathfrak{J}$ is (Poincaré) covariant if $\mathfrak{J}\big{(}W(g)\phi,\mathfrak{x}\big{)}=A\cdot\mathfrak{J}\big{(}\phi,g^{-1}\cdot\mathfrak{x}\big{)}$, $g=(\mathfrak{a},A)\in\tilde{\mathcal{P}}$ holds.
The introduction of a suitable dense space $\mathcal{D}$ in (5) takes account of the fact that $\mathfrak{J}$ may be given, as in the case of the massive scalar boson, as an integral operator, which is not defined for all wave functions $\phi$. Moreover, $\mathfrak{J}$ is supposed to be real. Actually, by checking the proof of (5), this is a consequence of $J_{0}$ being real and covariance.
(15) Lemma. A covariant current satisfies $J_{0}(\phi,\mathfrak{x})\geq|J(\phi,\mathfrak{x})|$ for $\mathfrak{x}\in\mathbb{R}^{4}$, $\phi\in\mathcal{D}$, if and only if $J_{0}(\phi;0,x)\geq 0$ for $x\in\mathbb{R}^{3}$, $\phi\in\mathcal{D}$.
Proof.
$J_{0}(\phi;0,x)\geq 0$ for $x\in\mathbb{R}^{3}$, $\phi\in\mathcal{D}$Leftrightarrow $\mathfrak{J}(W(g)^{-1}\phi;0,x)\cdot(1,0,0,0)\geq 0$ for $x\in\mathbb{R}^{3}$, $\phi\in\mathcal{D}$, $g=(\mathfrak{a},A)\in\tilde{\mathcal{P}}$Leftrightarrow $\mathfrak{J}(\phi;\mathfrak{a}+A\cdot(0,x))\cdot(A\cdot(1,0,0,0)\geq 0$ for $x\in\mathbb{R}^{3}$, $\phi\in\mathcal{D}$, $(\mathfrak{a},A)\in\tilde{\mathcal{P}}$Leftrightarrow $\mathfrak{J}(\phi;\mathfrak{x})\cdot\mathfrak{e}\geq 0$ for $\mathfrak{x}\in\mathbb{R}^{4}$, $\phi\in\mathcal{D}$, $\mathfrak{e}^{\cdot 2}=1$, $e_{0}>0$, whence the claim.∎
(16) Definition. A map $q:\mathcal{D}\to\mathbb{R}$ is said to be quadratic if (i) $q(\lambda\phi)=|\lambda|^{2}q(\phi)$ for $\lambda\in\mathbb{C}$, $\phi\in\mathcal{D}$ and (ii) $s_{q}(\phi,\phi^{\prime}):=\frac{1}{4}\sum_{\zeta=1,-1,\operatorname{i},-\operatorname{i}}\zeta q(\zeta\phi+\phi^{\prime})$ for $\phi,\phi^{\prime}\in\mathcal{D}$ is Hermitian sesquilinear. Note that $s_{q}(\phi,\phi)=q(\phi)$ holds by (i), whence (ii) is the polarization identity of $s_{q}$.
(17) Lemma. Assume (5.1) for $\phi\in\mathcal{D}$. Then, for every $x\in\mathbb{R}^{3}$, $J_{0}(\cdot\,;0,x)$ is quadratic.
Proof. Put $q_{x}:=J_{0}(\cdot\,;0,x)$. Note $\langle\phi,T(\Delta)\phi\rangle=\int_{\Delta}q_{x}(\phi)\operatorname{d}^{3}x$ and hence $\langle\phi,T(\Delta)\phi^{\prime}\rangle=\int_{\Delta}s_{q_{x}}(\phi,\phi^{\prime})\operatorname{d}^{3}x$, where $\langle\phi,T(\Delta)\phi^{\prime}\rangle$ is a Hermitian sesquilinear form.
       Hence, as to (5) (i), $\int_{\Delta}\big{(}q_{x}(\lambda\phi)-|\lambda|^{2}q_{x}(\phi)\big{)}\operatorname{d}^{3}x=0$ holds for every $\Delta$, whence the integrand is a.e. zero and by its continuity it is zero everywhere.
       Turn to (5) (ii). Show $s_{q_{x}}(\phi_{1}+\phi_{2},\phi^{\prime})-s_{q_{x}}(\phi_{1},\phi^{\prime})-s_{q_{x}}(\phi_{2},\phi^{\prime})=0$. Indeed, one has $\int_{\Delta}\big{(}s_{q_{x}}(\phi_{1}+\phi_{2},\phi^{\prime})-s_{q_{x}}(\phi_{1},\phi^{\prime})-s_{q_{x}}(\phi_{2},\phi^{\prime})\big{)}\operatorname{d}^{3}x=0$ for every $\Delta$, whence the claim. The remaining properties for $s_{q_{x}}$ regarding a Hermitian sesquilinear form hold analogously.∎
(18) Lemma. Let $\mathfrak{J}$ be covariant. Then $J_{0}(\cdot\,,\mathfrak{x})-J(\cdot\,,\mathfrak{x})\,e$ is quadratic on $\mathcal{D}$ for every $\mathfrak{x}\in\mathbb{R}^{4}$ and $e\in\mathbb{R}^{3}$, $|e|\leq 1$ if and only if $J_{0}(\cdot\,;0,x)$ is quadratic on $\mathcal{D}$ for every $x\in\mathbb{R}^{3}$.
Proof. $J_{0}(\phi;0,x)$ is quadratic regarding $\phi$ for every $x\in\mathbb{R}^{3}$Leftrightarrow $J_{0}(W(g)^{-1}\phi;0,x)$ is quadratic regarding $\phi$ for every $x\in\mathbb{R}^{3}$, $g=(\mathfrak{a},A)\in\tilde{\mathcal{P}}$Leftrightarrow $\mathfrak{J}(W(g)^{-1}\phi;0,x)\cdot(1,0,0,0)$ = $\mathfrak{J}(\phi;\mathfrak{a}+A\cdot(0,x))\cdot(A\cdot(1,0,0,0))$ is quadratic regarding $\phi$ for every $x\in\mathbb{R}^{3}$, $(\mathfrak{a},A)\in\tilde{\mathcal{P}}$Leftrightarrow $\mathfrak{I}(\phi;\mathfrak{x})\cdot\mathfrak{e}$ is quadratic regarding $\phi$ for every $\mathfrak{x}\in\mathbb{R}^{4}$, $\mathfrak{e}^{\cdot 2}=1$, $e_{0}>0$, whence the claim.∎
The main result follows.
(19) Theorem. Let $\mathfrak{J}$ be a conserved covariant current. Let $\phi\in\mathcal{D}$. Let $J_{0}$ satisfy
  • (a)
    $J_{0}(\phi;0,x)\geq 0$ for $x\in\mathbb{R}^{3}$, $\int J_{0}(\phi;0,x)\operatorname{d}^{3}x=||\phi||^{2}$, $J_{0}(\cdot\,;0,x)$ is quadratic for every $x\in\mathbb{R}^{3}$
  • (b)
    $J_{0}(\phi,\mathfrak{x})\leq C_{\phi}(1+x^{2})^{-2}$ for some $0<C_{\phi}<\infty$
Then there is a covariant AL $T$ satisfying for every achronal Borel set $\Delta$
$\langle\phi,T(\Delta)\phi\rangle=\int_{\varpi(\Delta)}\big{(}J_{0}(\phi;\tau(x),x)-J(\phi;\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x$ (1)
where $\tau:\varpi(\Delta)\to\mathbb{R}$ with $\operatorname{graph}\tau=\Delta$.
Proof.
Let $\phi\in\mathcal{D}$. By (a) and (5), $J_{0}(\phi,\mathfrak{x})\geq|J(\phi,\mathfrak{x})|$. Hence by (b) and (4), $\frac{1}{2}J_{0}(\phi,\mathfrak{x})+\frac{1}{2}C_{\phi}(1+x^{2})^{-2}\geq|J(\phi,\mathfrak{x})|$.
       Let $\Delta_{0}\in\mathcal{B}^{ach}$. There is a maximal achronal set $\Lambda\supset\Delta_{0}$ being the graph of the corresponding $1$-Lipschitz function $\tau:\mathbb{R}^{3}\to\mathbb{R}$. For Borel $\Delta\subset\Lambda$ put
$\pi_{\phi,\Lambda}(\Delta):=\int_{\varpi(\Delta)}\big{(}J_{0}(\phi;\tau(x),x)-J(\phi;\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x\>.$ (2)
As the integrand is nonnegative, $\pi_{\phi,\Lambda}$ is a $\sigma$-additive measure. By (a) and (4.3), $\pi_{\phi,\Lambda}(\Lambda)=||\phi||^{2}$. Due to (5), (5) $\phi\mapsto\pi_{\phi,\Lambda}(\Delta)$ is the quadratic form of a bounded Hermitian sesquilinear form on $\mathcal{D}$. Therefore by [11, Lemma 48] there is a bounded operator $T(\Delta)$, $0\leq T(\Delta)\leq I$ with $\pi_{\phi,\Lambda}(\Delta)=\langle\phi,T_{\Lambda}(\Delta)\phi\rangle$.
       Let $\phi\in\mathcal{H}$. By continuity on $\phi$, $\Delta\mapsto\langle\phi,T_{\Lambda}(\Delta)\phi\rangle$ is finitely additive. Actually it is $\sigma$-additive. Indeed, let $\Delta_{n}\downarrow_{n}\emptyset$ for Borel $\Delta_{n}\subset\Lambda$. Let $\epsilon>0$. Let $\phi^{\prime}\in\mathcal{D}$ with $||\phi-\phi^{\prime}||\leq\epsilon$. An obvious application of the triangle inequality yields $||\langle\phi,T_{\Lambda}(\Delta_{n})\phi\rangle-\langle\phi^{\prime},T_{\Lambda}(\Delta_{n})\phi^{\prime}\rangle||\leq||\phi-\phi^{\prime}||\,||\phi||+||\phi^{\prime}||\,||\phi-\phi^{\prime}||\leq 2\epsilon||\phi||+\epsilon^{2}\leq C\epsilon$ with $C<\infty$ independent of $\Delta_{n}$. As $\langle\phi^{\prime},T_{\Lambda}(\Delta_{n})\phi^{\prime}\rangle\to 0$, one infers $\langle\phi,T_{\Lambda}(\Delta_{n})\phi\rangle\to 0$, whence the claim.
       So $T_{\Lambda}$ is weakly $\sigma$-additive, which by [20, Theorem 4.28] implies the strong $\sigma$-additivity.
       Note that the definition of $T_{\Lambda}(\Delta_{0})$ via (2) does not depend on $\Lambda$. One may omit the index. Thus $T$ is an AL. It remains to show its covariance, which follows immediately by the subsequent proposition (5)(b).∎
(20) Remark. The assumption (5) (a) can be replaced by
  • (a’)
    $\int J_{0}(\phi;0,x)\operatorname{d}^{3}x=||\phi||^{2}$ and $\langle\phi,T(\Delta)\phi\rangle=\int_{\Delta}J_{0}(\phi;0,x)\operatorname{d}^{3}x$ with nonnegative operators $T(\Delta)$ for bounded Borel $\Delta\subset\mathbb{R}^{3}$
since, arguing as in (5), (a’) implies (a).
Regarding the notations see (5)(2).
(21) Proposition. Let $g=(\mathfrak{a},A)\in\tilde{\mathcal{P}}$. Then
       (a) the achronal set $g\cdot\Delta$ equals $\{(\tau_{g}(y),y):y\in\varpi(g\cdot\ \Delta)\}$ for $\tau_{g}(y):=(g\cdot(\tau(x),x))_{0}$ with $x:=S^{-1}(y)$, where $S:\mathbb{R}^{3}\to\mathbb{R}^{3}$, $S(x):=\varpi\big{(}g\cdot(\tau(x),x)\big{)}$ is a bijection.
       (b) $\pi_{W(g)^{-1}\phi,\Lambda}(\Delta)=\pi_{\phi,g\cdot\Lambda}(g\cdot\Delta)$.
Proof.
(a) Obviously $S$ is surjective. Let $S(x)=S(x^{\prime})$. Hence $\varpi\big{(}g\cdot(\tau(x)-\tau(x^{\prime}),x-x^{\prime})\big{)}=0$ with $\big{(}g\cdot(\tau(x)-\tau(x^{\prime}),x-x^{\prime})\big{)}^{\cdot 2}=\big{(}\tau(x)-\tau(x^{\prime}),x-x^{\prime}\big{)}^{\cdot 2}\leq 0$. Therefore also $(g\cdot(\tau(x)-\tau(x^{\prime}),x-x^{\prime}))_{0}=0$, whence $(g\cdot(\tau(x)-\tau(x^{\prime}),x-x^{\prime}))=0$. This means $(\tau(x)-\tau(x^{\prime}),x-x^{\prime})=0$. So $x=x^{\prime}$. In conclusion $S$ is bijective.
       Note $S(\varpi(\Delta))=\varpi(g\cdot\Delta)$. Therefore $\{(\tau_{g}(y),y):y\in\varpi(g\cdot\ \Delta)\}=\{\big{(}(g\cdot(\tau(x),x))_{0},S(x)\big{)}:x\in\varpi(\Delta)\}=\{g\cdot(\tau(x),x):x\in\varpi(\Delta)\}=g\cdot\Delta$.
       (b) By (5) (ii), $\pi_{W(g)^{-1}\phi,\Lambda}(\Delta)=\int_{\varpi(\Delta)}\mathfrak{J}\big{(}\phi,g\cdot(\tau(x),x)\big{)}\cdot\big{(}A\cdot(1,\nabla\tau(x))\big{)}\operatorname{d}^{3}x=\int_{\varpi(\Delta)}\mathfrak{J}\big{(}\phi,\tau_{g}(S(x)),S(x)\big{)}\cdot\big{(}A\cdot(1,\nabla\tau(x))\big{)}\operatorname{d}^{3}x$, which using the image of the Lebesgue measure $\lambda$ equals $\int_{S(\varpi(\Delta))}\mathfrak{J}\big{(}\phi,\tau_{g}(y),y\big{)}\cdot\big{(}A\cdot(1,\nabla\tau(S^{-1}(y))\big{)}\operatorname{d}S(\lambda)(y)$. Now recall $S(\varpi(\Delta))=\varpi(g\cdot\Delta)$ and note $\operatorname{d}S(\lambda)/\operatorname{d}\lambda=|\det D\,S^{-1}|=|\det D\,S(S^{-1}(\cdot))|^{-1}$.
       It remains to verify
$(1,\nabla\tau_{g}(y))=|\det D\,S(S^{-1}(y))|^{-1}\,A\cdot(1,\nabla\tau(S^{-1}(y)))$ (*)
which is easy in the case $A\in SU(2)$. So it suffices to check the case $g=\operatorname{e}^{\rho\sigma_{3}/2}$, $\rho\in\mathbb{R}$444 Explicitly $\operatorname{e}^{\,\rho\,\sigma_{3}/2}=\operatorname{diag}(\operatorname{e}^{\rho/2},\operatorname{e}^{-\rho/2})$ acts on $\mathbb{R}^{4}$ by $\left(\begin{array}[]{cccc}\cosh(\rho)&0&0&\sinh(\rho)\\ 0&1&0&0\\ 0&0&1&0\\ \sinh(\rho)&0&0&\cosh(\rho)\end{array}\right)$ for $\rho\in\mathbb{R}$ Put $c:=\cosh\rho$, $s:=\sinh\rho$, $z:=\nabla\tau(x)$, $x=S^{-1}(y)$. The rows of $\big{(}D\,S(x)\big{)}^{-1}$ are $(1,0,0)$, $(0,1,0)$, $\frac{1}{c+sz_{3}}(-sz_{1},-sz_{2},1)$. So the right side of (*) equals $\frac{1}{c+sz_{3}}(c+sz_{3},z_{1},z_{2},cz_{3}+s)$. On the left hand side $\nabla\tau_{g}(y))=\big{(}cz_{1}-\frac{(cz_{3}+s)sz_{1}}{c+sz_{3}},\dots,\frac{cz_{3}+s}{c+sz_{3}}\big{)}$. Hence (*) holds thus accomplishing the proof.∎
6 Covariant achronal localizations of the massive scalar boson
One recalls that the localizability of the massive scalar boson in Euclidean space is described by a Euclidean covariant normalized POVM $T$ on the Borel sets of $\mathbb{R}^{3}$, called a POL (Positive Operator Localization) in [5].
       To be explicit in the following we use the momentum representation with $L^{2}(\mathbb{R}^{3})$ being the space of states. By [5, (6.1), (11) Theorem] one has
  • $\langle\phi,T(\Delta)\phi\rangle=\int_{\Delta}J_{0}(\phi,x)\operatorname{d}^{3}x$
i.e., (5.1) holds, where the density of the probability of localization $J_{0}$ is given by
  • $J_{0}(\phi,x)=(2\pi)^{-3}\int\int\,\textsc{k}(k,p)\operatorname{e}^{\operatorname{i}(p-k)x}\overline{\phi(k)}\phi(p)\,\operatorname{d}^{3}k\,\operatorname{d}^{3}p$
for $\phi\in C_{c}$, i.e., continuous with compact support. Here k is any measurable normalized (i.e., $\textsc{k}(p,p)=1)$ rotational invariant positive definite separable kernel k on $\mathbb{R}^{3}\setminus\{0\}$.
       The aim is to extend $T$ to an AL which is Poincaré covariant under the representation $W$ describing the massive scalar boson [^5] Often one uses the antiunitarily equivalent $\operatorname{e}^{-\operatorname{i}\mathfrak{a}\cdot\,\mathfrak{p}}$.
  • $\big{(}W(\mathfrak{a},A)\phi\big{)}(p)=\sqrt{\epsilon(q)/\epsilon(p)}\,\operatorname{e}^{\operatorname{i}\mathfrak{a}\cdot\,\mathfrak{p}}\,\phi(q)$
    with $\epsilon(p):=\sqrt{m^{2}+p^{2}}$, $\mathfrak{p}:=(\epsilon(p),p)$, $\mathfrak{q}=(q_{0},q):=A^{-1}\cdot\mathfrak{p}$
following the considerations in sec. 5. Petzold and collaborators [13] show that $J_{0}$ is the zero component of a covariant conserved four-vector current $\mathfrak{J}:=(J_{0},J)$ if and only if
$\mathfrak{J}(\phi,\mathfrak{x})=(2\pi)^{-3}\int\int\,\mathfrak{K}(k,p)\operatorname{e}^{\operatorname{i}\big{(}(\epsilon(k)-\epsilon(p))x_{0}-(k-p)x\big{)}}\overline{\phi(k)}\phi(p)\,\operatorname{d}^{3}k\operatorname{d}^{3}p$ (6.1)
with $\phi\in C_{c}$. Here
$\mathfrak{K}(k,p):=\frac{(\epsilon(k)+\epsilon(p),k+p)}{2\sqrt{\epsilon(k)}\sqrt{\epsilon(p)}}g\big{(}\epsilon(k)\epsilon(p)-kp\big{)}$ (6.2)
where $g:[m^{2},\infty[\to\mathbb{R}$ is continuous with $g(m^{2})=1$ such that the zeroth component $K_{0}$ of $\mathfrak{K}$ is a positive definite kernel on $\mathbb{R}^{3}$ (see also [5, (55) Corollary]).
(22) Definition. $\mathfrak{K}$ in (6.2) is called a causal kernel if its zeroth component is positive definite on $\mathbb{R}^{3}$.
For a thorough analysis of the solutions $g$ see [5]. We mention $|g(t)|<g_{3/2}(t)$ if $t\neq m^{2}$, $g\neq g_{3/2}$, where $g_{r}(t):=(2m^{2})^{r}(m^{2}+t^{2})^{-r}$ for $r\geq 3/2$ denotes the basic series of solutions revealed by [13] and [14].
Henceforth we deal with the conserved covariant currents $\mathfrak{J}$ with causal kernel (6.1). For $\mathcal{D}:=C^{\infty}_{c}(\mathbb{R}^{3})$ the assumptions on $\mathfrak{J}$ in (5) are satisfied. $\mathfrak{J}(\phi,\cdot)$ is even smooth. Moreover $J_{0}$ satisfies (5.1) and hence (5)(a) by (5). Regarding the assumption (5)(b) one has
(23) Lemma. Let $\phi\in C^{\infty}_{c}(\mathbb{R}^{3})$. For $g$ in (6.2) assume $g\in C^{4}([m^{2},\infty[)$. Then (5)(b) holds.
Proof. Let $\alpha$ be a multiindex with $|\alpha|\leq 4$. Note $x^{\alpha}\operatorname{e}^{\operatorname{i}px}=(-\operatorname{i})^{|\alpha|}D_{p}^{\alpha}\operatorname{e}^{\operatorname{i}px}$. Multiply $J_{0}(\phi;0,x)$ in (6.1) by $x^{\alpha}$. After integration by parts the integrand is bounded by $\big{|}D_{p}^{\alpha}\big{(}K_{0}(k,p)\phi(p)\big{)}\big{|}\,|\phi(k)|$, which is integrable. So $|x^{\alpha}J_{0}(\phi;0,x)|$ is bounded by a constant independent of $x$.∎
       Note that $g$ from the basic series $g_{r}(t)=(2m^{2})^{r}(m^{2}+t^{2})^{-r}$, $r\geq 3/2$ is even $C^{\infty}$, in particular the distinguished $g=g_{3/2}$.
One summarizes (6.1), (6), (5).
(24) Theorem. Let $\mathfrak{J}$ be a covariant conserved current with causal kernel for the massive scalar boson. Assume $g\in C^{4}([m^{2},\infty[)$. Then there is an AL $T$ satisfying for every achronal Borel set $\Delta$ and $\phi\in C^{\infty}_{c}(\mathbb{R}^{3})$
$\langle\phi,T(\Delta)\phi\rangle=\int_{\varpi(\Delta)}\big{(}J_{0}(\phi;\tau(x),x)-J(\phi;\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x$
where $\tau:\varpi(\Delta)\to\mathbb{R}$ with $\operatorname{graph}\tau=\Delta$. $T$ is covariant.
We turn to the family of localizations of the massive scalar boson obtained out of its stress energy tensor [17]. For a thorough treatment see [11, sec. 6]. The family is indexed by the normalized future-directed timelike fourvectors $\mathfrak{n}$, i.e. $\mathfrak{n}^{\cdot 2}=1$, $n_{0}>0$. The related currents read still in the momentum representation with $\phi\in C_{c}^{\infty}(\mathbb{R}^{3})$
$\mathfrak{J}_{\mathfrak{n}}(\phi,\mathfrak{x})=(2\pi)^{-3}\int\int\,\mathfrak{K}_{\mathfrak{n}}(k,p)\operatorname{e}^{\operatorname{i}\big{(}(\epsilon(k)-\epsilon(p))x_{0}-(k-p)x\big{)}}\overline{\phi(k)}\phi(p)\,\operatorname{d}^{3}k\operatorname{d}^{3}p$ (6.3)
$\mathfrak{K}_{\mathfrak{n}}(k,p):=\frac{\mathfrak{k}\cdot\mathfrak{n}\;\mathfrak{p}+\mathfrak{p}\cdot\mathfrak{n}\;\mathfrak{k}-(m^{2}+\mathfrak{k}\cdot\mathfrak{p})\;\mathfrak{n}}{2\sqrt{\epsilon(k)}\sqrt{\epsilon(p)}}$ (6.4)
with $\mathfrak{p}:=(\epsilon(p),p)$, $\mathfrak{k}:=(\epsilon(k),k)$.
       One easily checks that $\mathfrak{J}_{\mathfrak{n}}(\phi,\cdot)$ is smooth bounded conserved and that $J_{\mathfrak{n},0}(\cdot\,;0,x)$ is quadratic for every $x\in\mathbb{R}^{3}$. Also one verifies the covariance
$\mathfrak{J}_{\mathfrak{n}}\big{(}W(g)\phi,\mathfrak{x}\big{)}=A\cdot\mathfrak{J}_{A^{-1}\cdot\mathfrak{n}}(\phi,g^{-1}\mathfrak{x})$ (6.5)
for all $\mathfrak{n}$, $g=(\mathfrak{a},A)$, $\mathfrak{x}$. In addition, for every $\mathfrak{n}$, $\phi$, one has $J_{\mathfrak{n},0}(\phi,;0,\cdot)\geq 0$ and $\int J_{\mathfrak{n},0}(\phi,;0,x)\operatorname{d}^{3}x=||\phi||^{2}$ as shown in [11, (64), Theorem 54]. Lemma (5), (5) hold for every $\mathfrak{J}_{\mathfrak{n}}$ by the same proofs due to (6.5). Finally, (5)(b) holds for every $J_{\mathfrak{n},0}$ by a proof analogous to that of (6). In summary, (5) applies to $\mathfrak{J}_{\mathfrak{n}}$. It follows
(25) Theorem. For every $\mathfrak{n}$ with $\mathfrak{n}^{\cdot 2}=1$, $n_{0}>0$, there is an AL $M^{\mathfrak{n}}$ satisfying for every achronal Borel set $\Delta$ and $\phi\in C^{\infty}_{c}(\mathbb{R}^{3})$
$\langle\phi,M^{\mathfrak{n}}(\Delta)\phi\rangle=\int_{\varpi(\Delta)}\big{(}J_{\mathfrak{n},0}(\phi;\tau(x),x)-J_{\mathfrak{n}}(\phi;\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x$
where $\tau:\varpi(\Delta)\to\mathbb{R}$ with $\operatorname{graph}\tau=\Delta$. One has the covariance $W(g)M^{\mathfrak{n}}(\Delta)W(g)^{-1}=M^{g\cdot\mathfrak{n}}(g\cdot\Delta)$.
7 Covariant representation of the causal logic for the massive scalar boson
The causal logic $\mathcal{C}$ is the lattice of Borel subsets of $\mathbb{R}^{4}$ which is partially ordered by set inclusion $\subset$ and which is generated and orthocomlemented by achronal separateness, i.e., the relation
$\mathfrak{x}\perp\mathfrak{y}\quad\Leftrightarrow\quad\mathfrak{x}\neq\mathfrak{y}\textnormal{ and }(\mathfrak{x}-\mathfrak{y})^{\cdot 2}\leq 0\>.$ (7.1)
The elements of $\mathcal{C}$ are called causally complete sets. In fact, they satisfy the requirement Δ=($\Delta=(\Delta^{\perp})^{\perp}$ where $M^{\perp}:=\{\mathfrak{x}\in\mathbb{R}^{4}\>:\>\mathfrak{x}\perp\mathfrak{y}\>,\forall\mathfrak{y}\in M\}$ is the causal complement of $M\subset\mathbb{R}^{4}$, and $(M^{\perp})^{\perp}$ is the causal completion of $M$.
       The causal logic is thoroughly studied by Cegla, Jadzyk in [8] and has been studied further on. The lattice $\mathcal{C}$ enjoys some appealing physical properties similar to the ones of the lattice of abstract elementary propositions of a quantum system [8] (see also [3]): it is possible to prove that it is $\sigma$-complete irreducible, orthomodular, atomic, atomistic exactly as for a quantum lattice, but fails to satisfy the covering law and separability.
       Since the work of [8] there persists the outstanding question how to construct covariant representations of the causal logic.
(26) Definition. Let $F(M)$ for $M\in\mathcal{C}$ be a bounded nonnegative operator on $\mathcal{H}$. Suppose $F(\emptyset)=0$, and $\sum_{n}F(M_{n})=I$ for every sequence $(M_{n})$ of mutually orthogonal sets in $\mathcal{C}$ such that $\bigvee_{n}M_{n}=\mathbb{R}^{4}$. Then the map $F$ is called a representation of the causal logic (RCL).
       Let $W$ be a unitary representation of $\tilde{\mathcal{P}}$. Then the RCL $F$ is said to be (Poincaré) covariant by means of $W$ if $F(g\cdot M)=W(g)F(M)W(g)^{-1}$ holds for $g\in\tilde{\mathcal{P}}$ and $M\in\mathcal{C}$.
The convergence of the above sum occurs in the weak operator topology (equivalently in the strong operator topology).
By the way a RCL $F$ is a map from the lattice $\mathcal{C}$ to the generalized $\sigma$ effect algebra of effects $E(\mathcal{H})$ on $\mathcal{H}$ [10]. This latter enjoys a weakened form of $\sigma$-complete orthocoplemented lattice structure. $F$ is a homomorphism of this weakened type of structure. In particular, it is easy to prove that a RCL is $\sigma$-additive, order-preserving and orthocomplement-preserving.
Only recently covariant RCL have been constructed. They concern quantum mechanical systems with definite spin $j\in\mathbb{N}_{0}/2$ and mass spectrum $\subset$ ]0,[]0$\infty$[] 0 , ∞ [ of positive Lebesgue measure [6]. Moreover, due to the results in [4], there is evidence that there are also RCL regarding the Dirac particles and the four Weyl fermions. These results have been obtainded by group theoretical methods.
       A different approach is to relate RCL to Poincaré covariant conserved (operator) density currents. See the works cited in [2]. A first concrete step in realizing a representation is done in [9] essentially showing (4.3) for a smooth conserved current with compact support.
       Recently an other idea of localization of the massive scalar boson in causally complete regions is pursued [18]. It uses the method of modular localization from AQFT. Given a Cauchy surface, via the modular localization map, to every state and to the causal completion of every its Borel subsets a probability of localization is attributed. This is asymptotically additive when increasing the distance between the regions of localization.
       There is the closed relationship (7) between AL and RCL, which is easy to expound. The set of determinacy of $M\subset\mathbb{R}^{4}$ is defined as
$M^{\sim}:=\{\mathfrak{x}:\forall\;\mathfrak{z}\textnormal{ with }\mathfrak{z}^{\cdot 2}>0\;\exists\,s\in\mathbb{R}\textnormal{ with }\mathfrak{x}+s\mathfrak{z}\in M\}\>.$
It consists of all points $\mathfrak{x}$ such that every timelike line through $\mathfrak{x}$ meets $M$. There is the remarkable result that, if $\Delta\in\mathcal{B}^{ach}$, then $\Delta^{\sim}=(\Delta^{\perp})^{\perp}$, whence $\Delta^{\sim}\in\mathcal{C}$ and conversely, if $M\in\mathcal{C}$ then $M=\Delta^{\sim}$ for $\Delta\in\mathcal{B}^{ach}$ maximal achronal in $M$. For details see [8], [4], [6]. This feature of spacetime is the reason for the following one-to-one correspondence.
(27) Proposition. See [6, (19), (20)].
       (a) Let $F$ be an RCL. Set $T(\Delta):=F(\Delta^{\sim})$ for $\Delta\in\mathcal{B}^{ach}$. Then $T$ is an AL. If $F$ is covariant by means of $W$, then so is $T$.
       (b) Let $T$ be an AL. Then there is a unique RCL $F$ with $F(\Delta^{\sim})=T(\Delta)$ for $\Delta\in\mathcal{B}^{ach}$. If $T$ is covariant by means of $W$, then so is $F$.
It suffices to join up (6), (7).
(28) Theorem. To every causal kernel $\mathfrak{K}$ with $g\in C^{4}([m^{2},\infty[)$ there is a unique covariant RCL $F$ for the massive scalar boson such that
$\langle\phi,F(\Delta^{\sim})\phi\rangle=\langle\phi,T(\Delta)\phi\rangle$
holds for $\phi\in C^{\infty}_{c}(\mathbb{R}^{3})$ and every achronal Borel set $\Delta$. Here $\langle\phi,T(\Delta)\phi\rangle$ is given in (6) with $\mathfrak{J}$ from (6.1).
Thus, apparently for the first time, a covariant RCL for a quantum mechanical system with definite mass is achieved.
       Similarly one obtains from (6) a covariant family of RCL related to the stress energy tensor of the massive scalar boson.
(29) Theorem. For every $\mathfrak{n}$ with $\mathfrak{n}^{\cdot 2}=1$, $n_{0}>0$, there is a unique RCL $F^{\mathfrak{n}}$ for the massive scalar boson such that
$\langle\phi,F^{\mathfrak{n}}(\Delta^{\sim})\phi\rangle=\langle\phi,M^{\mathfrak{n}}(\Delta)\phi\rangle$
holds for $\phi\in C^{\infty}_{c}(\mathbb{R}^{3})$ and every achronal Borel set $\Delta$. Here $\langle\phi,M^{\mathfrak{n}}(\Delta)\phi\rangle$ is given in (6). One has the covariance $W(g)F^{\mathfrak{n}}(\Delta)W(g)^{-1}=F^{g\cdot\mathfrak{n}}(g\cdot\Delta)$.
Acknowledgements
V.M. and C. De R. are grateful to Silvano Delladio and Andrea Marchese for useful suggestions and discussions, and acknowledge that this work has been written within the activities of INdAM-GNFM.