### v4

# Achronal localization and representation of the causal logic from a conserved current, application to the massive scalar boson

Abstract
Covariant achronal localizations are gained out of covariant conserved currents computing their flux passing through achronal surfaces. This general method applies to the probability density currents with causal kernel regarding the massive scalar boson. Due to the one-to-one correspondence between (covariant) achronal localizations and (covariant) representations of the causal logic thus, apparently for the first time, a covariant representation of the causal logic for an elementary relativistic quantum mechanical system has been achieved. Similarly one derives the covariant family of representations of the causal logic related to the stress energy tensor of the massive scalar boson.
While reaching this result the divergence theorem is proven for open sets with almost Lipschitz boundary.
1 Notations and notions
Minkoski spacetime. Vectors in $\mathbb{R}^{4}$ are denoted by $\mathfrak{x}=(x_{0},x)$ with $x:=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$. Let $\varpi:\mathbb{R}^{4}\to\mathbb{R}^{3}$ denote the projection $\varpi(\mathfrak{x}):=x$. Representing Minkowski spacetime by $\mathbb{R}^{4}$ the Minkowski product of $\mathfrak{a},\mathfrak{a}^{\prime}\in\mathbb{R}^{4}$ is given by $\mathfrak{a}\cdot\mathfrak{a}^{\prime}:=a_{0}a_{0}^{\prime}-aa^{\prime}$, where for vectors $a,a^{\prime}$ in $\mathbb{R}^{3}$ the scalar product $a_{1}a^{\prime}_{1}+a_{2}a^{\prime}_{2}+a_{3}a^{\prime}_{3}$ is denoted by $aa^{\prime}$. Often we use the notation $\mathfrak{a}^{\cdot 2}:=\mathfrak{a}\cdot\mathfrak{a}$.
Poincaré group. $\tilde{\mathcal{P}}=ISL(2,\mathbb{C})$ is the universal covering group of the proper orthocronous Poincaré group. It acts on $\mathbb{R}^{4}$ as
$g\cdot\mathfrak{x}:=\mathfrak{a}+\varLambda(A)\mathfrak{x}\quad\text{ for }g=(\mathfrak{a},A)\in\tilde{\mathcal{P}},\,\mathfrak{x}\in\mathbb{R}^{4}$ (1.1)
where $\varLambda:SL(2,\mathbb{C})\to O(1,3)_{0}$ is the universal covering homomorphism onto the proper orthochronous Lorentz group. For short one writes $A\equiv(0,A)$, $\mathfrak{a}\equiv(\mathfrak{a},I_{2})$, and $A\cdot\mathfrak{x}=\varLambda(A)\mathfrak{x}$. For $M\subset\mathbb{R}^{4}$ and $g\in\tilde{\mathcal{P}}$ define $g\cdot M:=\{g\cdot\mathfrak{x}:\mathfrak{x}\in M\}$.
       The group operation on $\tilde{\mathcal{P}}$ reads $(\mathfrak{a},A)(\mathfrak{a}^{\prime},A^{\prime})=(\mathfrak{a}+A\cdot\mathfrak{a}^{\prime},AA^{\prime})$ with identity element $(0,I_{2})$ and inverse $(\mathfrak{a},A)^{-1}=(-A^{-1}\cdot\mathfrak{a},A^{-1})$.
Spacetime relations. The fourvector $\mathfrak{z}\in\mathbb{R}^{4}\setminus\{0\}$ is called timelike, lightlike, causal if $|z_{0}|>|z|$, $|z_{0}|=|z|$, $|z_{0}|\geq|z|$, respectively. It is future-directed if $z_{0}>0$.
       The set $\mathfrak{a}+\mathbb{R}\mathfrak{z}$ for $\mathfrak{a},\mathfrak{z}\in\mathbb{R}^{4}$, $\mathfrak{z}\neq 0$ is called a line. The line is timelike, lightlike, causal, if so is $\mathfrak{z}$.
       A set $\Delta\subset\mathbb{R}^{4}$ is said to be spacelike if different points of $\Delta$ are spacelike separated, i.e., $|x_{0}-y_{0}|<|x-y|$ for $\mathfrak{x},\mathfrak{y}\in\Delta$, $\mathfrak{x}\neq\mathfrak{y}$.
Achronal sets. A set $\Delta\subset\mathbb{R}^{4}$ is said to be achronal if the points of $\Delta$ are achronal separated, i.e., $|x_{0}-y_{0}|\leq|x-y|$ for $\mathfrak{x},\mathfrak{y}\in\Delta$.
       Clearly every spacelike set is achronal. By definition $\Delta$ is maximal achronal if $\Delta$ is not properly contained in an achronal set. An achronal set is maximal achronal if and only if it meets every timelike line.
       Cauchy surfaces in Minkowski space are maximal achronal sets, which meet all lightlike lines. The converse holds also true [8, Moretti (36)]. The mass shell is an example of a spacelike maximal achronal set, which is not a Cauchy surface. Being maximal it is not even contained in a Cauchy surface.
       Note also that every achronal set is contained in a maximal one and that a maximal achronal set is always closed.
       Most important is the fact that a set $\Delta\subset\mathbb{R}^{4}$ is achronal if and only if it is the graph of a $1$-Lipschitz function $\tau:\varpi(\Delta)\to\mathbb{R}$. It is maximal achronal iff $\varpi(\Delta)=\mathbb{R}^{3}$.
       Let the maximal achronal set $\Lambda=\{\big{(}\tau(x),x\big{)}:x\in\mathbb{R}^{3}\}$ be endowed with the induced topology from $\mathbb{R}^{4}$. Then $j:\mathbb{R}^{3}\to\Lambda$, $j(x):=(\tau(x),x)$ is a homeomorphism, $\{\mathfrak{x}\in\mathbb{R}^{4}:x_{0}<\tau(x)\}$ and $\{\mathfrak{x}\in\mathbb{R}^{4}:x_{0}>\tau(x)\}$ are open with boundary $\Lambda$, and $\Lambda$ is path-connected. In particular $\Lambda$ is a topological hypersurface. Clearly $\Lambda\subset\mathbb{R}^{4}$ has no interior points.
       For more details on achronal sets see [8, sec. 2.2].
2 Introduction
Achronal localization. There are plain physical grounds why in order to comply with causality, localization of a relativistic quantum mechanical system has to occur in all achronal regions of spacetime. This means that to every achronal Borel set $\Delta\subset\mathbb{R}^{4}$ (see sec. 1, Achronal sets) there is assigned an effect: a nonnegative operator $T(\Delta)$, bounded by $I$, acting on the Hilbert space of states. The expectation value $\langle\phi,T(\Delta)\phi\rangle$ is the probability of localization in $\Delta$ of the system in the state $\phi$. As argued further in [8] the probabilities of localization regarding achronally separated regions countably add up and yield $1$ for every maximal achronal region. In summary $T$ is a normalized positive operator valued measure (POVM) on every maximal achronal set. Moreover, by relativistic symmetry, $T$ is Poincaré covariant with respect to the unitary representation of the Poincaré group corresponding to kinematical transformations of the states.
       Achronal localization is causal. It satisfies the so-called Einstein causality in the most direct interpretation, which requires from localization that the probability of localization in a region of influence cannot be less than that in the region of actual localization. Details about this causality condition CC and achronal localization are furnished in sec. 5.
       Heuristically one may imagine a massive free particle to be represented by a timelike straight world line. The particle is considered to be localized in a spacetime region if its world line crosses the region. This simple heuristic picture of localizability implies rather stringent achronal localization [8, sec. 3.3]. Since the maximal achronal sets are those sets, which are met by every timelike straight line just once, one infers that a massive particle is localized in every maximal achronal region. Hence normalization follows. The present results on the localization of the massive scalar boson evidence the usefulness of this heuristical reasoning.
       The question is how to realize an apparatus, which ascertains the presence of a particle in an achronal spacetime region $\Delta$. From a physical point of view it is reasonable to assume the region to be spacelike piecewise flat. Presume that localization in $3D$ Euclidean space is feasible. This implies by relativistic symmetry that for every flat piece $\Gamma$ there is an apparatus suited to ascertain the presence of the particle in that region $\Gamma$. Being spacelike separated all these apparatus constitute one apparatus able to ascertain the presence of the particle in $\Delta$.
       Localization on spacelike Cauchy surfaces. Just in order to meet the requirements of causal localization [13] extended localization from flat spacelike regions, as up to then commonly considered, to proper spacetime regions studying thoroughly POVM on spacelike smooth Cauchy surfaces (which are a special case of maximal achronal sets). The POVM are coherent in the sense that the localization operators assigned to a region contained in the intersection of two different spacelike smooth Cauchy surfaces coincide. An important result is that the localization considered in [13] satisfies a rather general causality requirement. The results in [13] were achieved by a suitable use of the divergence theorem for volume forms. It was done by taking advantage of relatively recent advanced results in Lorentzian geometry concerning the extension of acausal manifolds with boundary to spacelike smooth Cauchy surfaces. These achievements actually straightforwardly extend to spacelike $C^{1}$ Cauchy surfaces.
Here we do the final step towards a fully-fledged achronal localization notion as described above.
       The construction of an achronal localization goes along the lines of the construction of the Cauchy localization in [13]. However there are some technical difficulties, which arise due to the facts (i) that a maximal achronal set in general is not a Cauchy surface as it is not met by every lightlike line and (ii) that an achronal set in general is not $C^{1}$, which means that it is not the graph of a $C^{1}$ function. Achronal sets are representable by Lipschitz functions, this only regularity condition should be therefore assumed.
       Divergence theorem. The divergence theorem plays a crucial role in the construction of concrete achronal localization notions. In order to overcome the difficulties pointed out above, we prove the divergence theorem for open bounded subsets of $\mathbb{R}^{n}$ with almost Lipschitz boundary such that the boundary has finite $(n-1)$-dimensional Hausdorff measure and the irregular points of the boundary are contained in a compact set of zero $(n-1)$-dimensional Minkowski content. This extension (3) of the divergence theorem is new. In view of applications we add some remarks about how to verify the assumptions of the theorem. At the beginning of sec. 4 it is expounded how (3) applies in showing the crucial technical result (4).
       Flux through maximal achronal sets. Using the prior result the crucial result is proven that the future-directed flux passing through a spacelike Cauchy surface of a conserved bounded zero or causal future-directed $C^{1}$-current is the same for all these surfaces (4).
       The result of (4) is extended to all maximal achronal sets containing the origin under the assumption on the decay (5)(b) of the probability current. Due to this decay the extension is achieved simply by flatten the maximal achronal sets becoming $\gamma$-achronal for $\gamma<1$. The assumed decay is determined by the free relativistic time evolution of a massive particle. Regarding the application to the massive scalar boson this assumption turns out to be a technicality satisfied for all relevant currents.
       Construction of achronal localization. At this juncture one is ready to derive the main result (5). Roughly speaking, every covariant conserved $C^{1}$ current with bounded zeroth component being positive quadratic on Euclidean space determines by the flux through the achronal sets a covariant achronal localization.
       Application to the massive scalar boson. The localizations of the elementary particle being the massive scalar boson in Euclidean space (i.e., the Euclidean covariant positive operator valued normalized measures in $\mathbb{R}^{3}$, called POL in [8]) are determined by the integrals of a positive definite rotational invariant probability density $J_{0}$ over the regions of localization [8, sec. 6]. $J_{0}$ is the zeroth component of a covariant conserved current $\mathfrak{J}$ if and and only if the kernel is causal (6). This result is by [15]. Causal kernels have been studied in [15], [19], [8]. Under the physically irrelevant condition that the causal kernel is $C^{4}$ one verifies that $\mathfrak{J}$ satisfies the assumption in (5) thus giving rise to a covariant achronal localization of the massive scalar boson.
       Analogously one obtains a covariant family of achronal localizations related to the stress energy tensor of the massive scalar boson [23].
       Representation of the causal logic. Every (covariant) achronal localization determines uniquely a (covariant) representation of the causal logic and vice versa [8, (21)]. Obviously this one-to-one correspondence emphasizes further the relevance of achronal localization.
       Hence for the first time a covariant representation of the causal logic for an elementary relativistic quantum mechanical system is achieved. Also a covariant family of representations of the causal logic is derived corresponding to the above mentioned family of achronal localizations related to the stress energy tensor.
3 Divergence theorem on open sets with almost Lipschitz boundary
The following is inspired by Maggi [21, Remark 9.5, Theorem 9.6]. Repeatedly one refers to the well-known Area Formula in geometric measure theory first proved by H. Federer, see [14, 3.2., in particular 3.2.20], citing the easier accessible [21, Theorem 9.1]. Part (a) of the proof of the preparatory result (3) is the standard step familiar from the proof of the classical Gauß’s theorem. See e.g. [1, proof of (A8-16)] or the more elementary Step one of the proof of [21, Theorem 9.3].
As to the notation, for $x\in\mathbb{R}^{n}$, $z\in\mathbb{R}^{n-1}$, $s>0$ put $x^{\prime}:=(x_{1},\dots,x_{n-1})$ and $x=(x^{\prime},x_{n})$, and let $B_{s}(x):=\{y\in\mathbb{R}^{n}:|y-x|<s\}$, $C(x,s):=\{y\in\mathbb{R}^{n}:|y^{\prime}-x^{\prime}|<s,|y_{n}-x_{n}|<s\}$, $D(z,s):=\{y\in\mathbb{R}^{n-1}:|y-z|<s\}$. Given an outer measure $\omega$ on $\mathbb{R}^{n}$ and $X\subset\mathbb{R}^{n}$, the trace (restriction) of $\omega$ on $X$ is the outer measure $\omega|X$ on $\mathbb{R}^{n}$ given by $\omega|X(F):=\omega(F\cap X)$. $\mathcal{H}^{n-1}$ denotes the $n-1$-dimensional Hausdorff measure on $\mathbb{R}^{n}$.
(1) Definition. Let $E\subset\mathbb{R}^{n}$ be open. $E$ has almost Lipschitz boundary $\partial E$ if there is a localization of $\partial E$ as follows. There is a closed set $M_{0}\subset\partial E$ with $\mathcal{H}^{n-1}(M_{0})=0$ such that for every $x\in M:=\partial E\setminus M_{0}$ there exists $s\in]0,\infty[$ with, up to rotation [^1] This means that there exists an orthonormal basis $e_{1},\dots,e_{n}$ of $\mathbb{R}^{n}$ such that (a),(b) hold for the coordinates of $x$ with respect to this basis.,
  • (a)
    $C(x,s)\cap E=\{y\in C(x,s):y_{n}>u(y^{\prime})\}$
  • (b)
    $C(x,s)\cap\partial E=C(x,s)\cap M=\{y\in C(x,s):y_{n}=u(y^{\prime})\}$
for some Lipschitz function $u:D(x^{\prime},s)\to\mathbb{R}$. Being $G$ the set of points of differentiability of $u$ define the unit vector field
$\nu:\operatorname{graph}u|_{G}\to\mathbb{R}^{n},\quad\nu(z,u(z)):=\frac{(\nabla u(z),-1)}{|(\nabla u(z),-1)|}$
Recall that by Rademacher’s theorem $G$ is the complement of a Lebesgue null set and that $\nabla u$ is measurable.
Henceforth $E\subset\mathbb{R}^{n}$ is open with almost Lipschitz boundary. The notation refers to (3).
(2) Lemma. $\mathcal{H}^{n-1}|C(x,s)\cap M$ is a Radon measure and $\mathcal{H}^{n-1}\big{(}(C(x,s)\cap M)\setminus\operatorname{graph}u|_{G}\big{)}=0$.
Proof.
Put $D:=D(x^{\prime},s)\cap\{z\in\mathbb{R}^{n-1}:|u(z)-x_{n}|<s\}$. Note $\operatorname{graph}\,u|_{D}=C(x,s)\cap\operatorname{graph}\,u=C(x,s)\cap\partial E=C(x,s)\cap M$.
       $\mathcal{H}^{n-1}|C(x,s)\cap M=\mathcal{H}^{n-1}|\operatorname{graph}\,u|_{D}$ is a Radon measure due to the area formula [21, Theorem 9.1] . Moreover, one has $\mathcal{H}^{n-1}(\operatorname{graph}\,u|_{D(x^{\prime},s)\setminus G})=\int_{D(x^{\prime},s)\setminus G}|(\nabla u(z),-1)|\operatorname{d}z=0$. ∎
The following preparatory result (3) comprises by the case $M_{0}=\emptyset$ the divergence theorem on bounded open sets with Lipschitz boundary (see also e.g. [21, Remark 9.5] and [1, A8.8]).
(3) Proposition. Let $E$ be bounded. Let $\varphi\in C_{c}^{1}(\mathbb{R}^{n})$ vanish on a neighborhood of $M_{0}$. Then
$\int_{E}\nabla\varphi\;\operatorname{d}\mathcal{L}^{n}=\int_{\partial E}\varphi\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}$
holds. Here $\nu_{E}$ is a $\mathcal{H}^{n-1}$-a.e. determined unit vector field on $\partial E$. Locally $\nu_{E}$ coincides with $\nu$ in (3).
Proof. (a) Assume first $\varphi\in C_{c}^{1}(C(x,s))$. By (3) the surface integral is well-defined. Following step one of the proof of [21, Theorem 9.3] the result follows due to $C(x,s)\cap\partial E=C(x,s)\cap M$ by definition (3)(b). Note that the weak gradient $\nabla f_{\delta}$ equals $\frac{1}{2\delta}(-\nabla u(z),1)$ at $z\in G$.
       (b) Now the vector field $\nu_{E}$ is constructed. Applying the result in (a) to all $\varphi\in C_{c}^{1}(C(x,s)\cap C(\bar{x},\bar{s}))$ shows that $\nu$ and $\bar{\nu}$ coincide a.e. on their common domain. There are countably many $C(x_{m},s_{m})$ from the localization in (3), which cover $M$. (Indeed, $M_{0}$ is closed and hence $G_{\delta}$. Since $\partial E$ is compact it follows that $M=\partial E\setminus M_{0}$ is $\sigma$-compact.) Thus we may compose from the corresponding $\nu_{m}$ by means of [21, Theorem 9.1] an $\mathcal{H}^{n-1}$-a.e. determined unit vector field $\nu_{E}$ on $\partial E$ which locally coincides a.e. with $\nu$ from (3).
       (c) As to the general case let $A\supset M_{0}$ be open with $\varphi|_{A}=0$. Note $\overline{E}\setminus A=(E\cup M)\setminus A\subset E\cup\bigcup_{x\in M}C(x,s)$. Since $\overline{E}\setminus A$ is compact it is covered by $E$ and finitely many $C(x,s)$. Hence one obtains a finite open cover of $\overline{E}$ by $U_{0}:=E$, $U_{k}$ being some $C(x,s)$ for $k=1,\dots,N$, and $U_{N+1}:=A$.
       Let $(\eta_{k})_{k=0,\dots,N+1}$ be a partition of unity for $\overline{E}$ subject to this cover [1, 4.20], i.e. $\eta_{k}\in C^{\infty}_{c}(\mathbb{R}^{n})$, $\operatorname{supp}\eta_{k}\subset U_{k}$, $\eta_{k}\geq 0$, and $\sum_{k=0}^{N+1}\eta_{k}(x)=1$ for $x\in\overline{E}$. Then
  • $\int_{E}\nabla(\eta_{0}\varphi)\operatorname{d}\mathcal{L}^{n}=0$ by the fundamental theorem of calculus, and obviously $\int_{\partial E}(\eta_{0}\varphi)\nu_{E}\operatorname{d}\mathcal{H}^{n-1}=0$
  • $1\leq k\leq N$: $\int_{E}\nabla(\eta_{k}\varphi)\operatorname{d}\mathcal{L}^{n}=\int_{\partial E}(\eta_{k}\varphi)\nu_{E}\operatorname{d}\mathcal{H}^{n-1}$ by (a), (b)
  • $k=N+1$: $\int_{E}\nabla(\eta_{N+1}\varphi)\operatorname{d}\mathcal{L}^{n}=0$ and $\int_{\partial E}(\eta_{N+1}\varphi)\nu_{E}\operatorname{d}\mathcal{H}^{n-1}=0$ as $\varphi|_{A}=0$
whence $\int_{E}\nabla\varphi\operatorname{d}\mathcal{L}^{n}=\sum_{k=1}^{N}\int_{E}\nabla(\eta_{k}\varphi)\operatorname{d}\mathcal{L}^{n}=\sum_{k=1}^{N}\int_{\partial E}(\eta_{k}\varphi)\nu_{E}\operatorname{d}\mathcal{H}^{n-1}=\int_{\partial E}\varphi\nu_{E}\operatorname{d}\mathcal{H}^{n-1}$.
       (d) In conclusion consider a further localization of $\partial E$ according (3) with $\tilde{\nu}_{E}$ the related unit vector field by (b). Then $\int_{\partial E}\varphi\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}=\int_{\partial E}\varphi\,\tilde{\nu}_{E}\;\operatorname{d}\mathcal{H}^{n-1}$ holds for all $\varphi\in C_{c}^{1}(\mathbb{R}^{n})$ with $\operatorname{supp}\varphi\cap M_{0}=\emptyset$, whence $\nu_{E}=\tilde{\nu}_{E}$ $\mathcal{H}^{n-1}$-a.e.∎
(4) Lemma. Let $X\subset\mathbb{R}^{n}$ be Borel. Then $\mathcal{H}^{n-1}|X$ is Borel regular.
Proof. Let $F\subset\mathbb{R}^{n}$. Since $\mathcal{H}^{n-1}$ is Borel regular, there are Borel sets $A,B\subset\mathbb{R}^{n}$ with $F\cap X\subset A$, $F\setminus X\subset B$ and $\mathcal{H}^{n-1}(F\cap X)=\mathcal{H}^{n-1}(A)$, $\mathcal{H}^{n-1}(F\setminus X)=\mathcal{H}^{n-1}(B)$. Then the Borel set $C:=(A\cap X)\cup(B\setminus X)$ satisfies $F\subset C$ with $F\cap X=A\cap X=C\cap X$ so that $\mathcal{H}^{n-1}(F\cap X)=\mathcal{H}^{n-1}(C\cap X)$, i.e., $\mathcal{H}^{n-1}|X(F)=\mathcal{H}^{n-1}|X(C)$. ∎
We are in a position to state and prove our version (3) of the Divergence Theorem. It distinguishes itself in the weakest hypotheses concerning $\partial E$ regularity we know. Thus (3) applies to the proof of the crucial technical result (4).
(5) Theorem. Let $E$ be open bounded with almost Lipschitz boundary. Assume $\mathcal{H}^{n-1}(\partial E)<\infty$ and the Minkowski content $\mathcal{M}^{n-1}(M_{0})=0$. Then $\mathcal{H}^{n-1}|\partial E$ is a Radon measure and for $\varphi\in C_{c}^{1}(\mathbb{R}^{n})$
$\int_{E}\nabla\varphi\;\operatorname{d}\mathcal{L}^{n}=\int_{\partial E}\varphi\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}$
holds. Here $E$ can be replaced by $\overline{E}$ as $\mathcal{L}^{n}(\partial E)=0$, and $\partial E$ can be replaced by $M$ as $\mathcal{H}^{n-1}(M_{0})=0$.
Proof.
$\mathcal{H}^{n-1}|\partial E$ is a Radon measure by (3). Let $\delta>0$ and let $A_{\delta}$ denote the $\delta$-neighborhood of $M_{0}$. Then
  • $\mathcal{H}^{n-1}(\partial E\cap A_{\delta})\to 0$ for $\delta\to 0$
  • $\delta^{-1}\mathcal{L}^{n}(A_{\delta})\to 0$ for $\delta\to 0$; in particular, $\mathcal{L}^{n}(A_{\delta})\to 0$
Indeed, the first claim holds as $\mathcal{H}^{n-1}(\partial E\cap A_{\delta})<\infty$ and $A_{\delta}\downarrow_{\delta}M_{0}$. The very definition of the Minkowski content [14, 3.2.27] implies the second claim.
       Furthermore by [1, 4.19] there is $f\in C^{\infty}_{c}(\mathbb{R}^{n})$ with $\operatorname{supp}f\subset A_{\delta}$, $0\leq f\leq 1$, $f|_{M_{0}}=1$, and $|\nabla f|\leq C/\delta$, where the finite constant $C$ does not depend on $\delta$.
       Note that (3) applies to $(1-f)\varphi$. So $\int_{E}\nabla\varphi\operatorname{d}\mathcal{L}^{n}=\int_{\partial E}(1-f)\varphi\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}+\int_{E}\nabla(f\varphi)\operatorname{d}\mathcal{L}^{n}=\int_{\partial E}\varphi\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}-\int_{\partial E}f\varphi\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}+\int_{E}(\nabla f)\varphi\operatorname{d}\mathcal{L}^{n}+\int_{E}f\nabla\varphi\operatorname{d}\mathcal{L}^{n}$. The last three summands vanish as $\delta\to 0$.
       Indeed, $|\int_{\partial E}f\varphi\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}|\leq\int_{\partial E}1_{A_{\delta}}|\varphi|\operatorname{d}\mathcal{H}^{n-1}\leq||\varphi||_{\infty}\mathcal{H}^{n-1}(\partial E\cap A_{\delta})\to 0$. Next, the crucial one, $|\int_{E}(\nabla f)\varphi\operatorname{d}\mathcal{L}^{n}|\leq\int_{E}C\delta^{-1}1_{A_{\delta}}|\varphi|\operatorname{d}\mathcal{L}^{n}=||\varphi||_{\infty}C\delta^{-1}\mathcal{L}^{n}(A_{\delta})\to 0$. Finally one has $|\int_{E}f\nabla\varphi\operatorname{d}\mathcal{L}^{n}|\leq\int_{E}1_{A_{\delta}}||\nabla\varphi||\operatorname{d}\mathcal{L}^{n}\leq||\nabla\varphi||_{\infty}\mathcal{L}^{n}(A_{\delta})\to 0$, thus accomplishing the proof.∎
An obviously equivalent formulation of (3) is
(6) Corollary. Let $E$ be as in (3) and $\mathfrak{v}\in C_{c}^{1}(\mathbb{R}^{n},\mathbb{R}^{n})$ be a vector field. Then
$\int_{E}\operatorname{div}\,\mathfrak{v}\;\operatorname{d}\mathcal{L}^{n}=\int_{\partial E}\mathfrak{v}\,\nu_{E}\;\operatorname{d}\mathcal{H}^{n-1}$
where $\operatorname{div}\,\mathfrak{v}:=\sum_{i=1}^{n}\partial_{i}v_{i}$ denotes the divergence of $\mathfrak{v}$ and $\mathfrak{v}\,\nu_{E}$ is the $\mathbb{R}^{n}$-scalar product of $\mathfrak{v}$ and $\nu_{E}$.
In view of an application of (3), (3) the following remarks on $\mathcal{H}^{n-1}(\partial E)<\infty$ and $\mathcal{M}^{n-1}(M_{0})=0$ may be useful. Note first that $\mathcal{M}^{n-1}(M_{0})=0$ implies $\mathcal{H}^{n-1}(M_{0})=0$.
       According to [21, Theorem 8.1], $\mathcal{H}^{n-1}(D)<\infty$ if $D$ is the image under an injective Lipschitz function on $\mathbb{R}^{n-1}$ in $\mathbb{R}^{n}$ of a Lebesgue measurable set $C$ of finite Lebesgue measure. If $C$ is a Lebesgue null set then $\mathcal{H}^{n-1}(D)=0$.
(7) Corollary. Let $\partial E$ be covered by finitely many sets $D_{i}$, each $D_{i}$ being the image under an injective Lipschitz function on $\mathbb{R}^{n-1}$ in $\mathbb{R}^{n}$ of a Lebesgue measurable $C_{i}$ of finite Lebesgue measure. Then $\mathcal{H}^{n-1}(\partial E)<\infty$.
If the images $D_{i}$ of countably many Lebesgue null sets $C_{i}$ cover $M_{0}$, then $\mathcal{H}^{n-1}(M_{0})=0$.
       According to [14, Theorem 3.2.39], $\mathcal{H}^{n-1}(D)=\mathcal{M}^{n-1}(D)=0$ holds, if $D\subset\mathbb{R}^{n}$ is closed and if there is a Lipschitz function $h:\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ mapping some bounded $C\subset\mathbb{R}^{n-1}$ onto $D$ or, equivalently, if $h$ is locally Lipschitz defined on $\overline{C}$ with $h(C)=D$. The latter holds because $\overline{C}$ is compact.
(8) Proposition. Let $M_{0}$ be the union of finitely many sets $D_{i}$ of Minkowski content $\mathcal{M}^{n-1}(D_{i})=0$. Then $\mathcal{M}^{n-1}(M_{0})=0$.
Moreover, $\mathcal{M}^{n-1}(D_{i})=0$ holds if $\mathcal{H}^{n-1}(D_{i})=0$ and if $D_{i}$ is the image of a compact set under a locally Lipschitz function.
Proof.
It remains to prove the first part of the assertion. The finite subadditivity of the upper Minkowski content yields $\mathcal{M}^{*n-1}(M_{0})\leq\sum_{i}\mathcal{M}^{*n-1}(D_{i})=\sum_{i}\mathcal{M}^{n-1}(D_{i})=0$ implying $\mathcal{M}^{n-1}(M_{0})=0$. ∎
4 Flux passing through a maximal achronal set
The result (4) essentially is an extension of [13, Proposition 37] to spacelike Cauchy surfaces in Minkowski spacetime. Note that [13, Proposition 37] is equally valid for a $C^{1}$ vector field in place of a smooth one. General spacetime theories study Cauchy surfaces. These are the sets which meet every inextendible timelike smooth curve exactly once [25, Chapter 14, Definition 28]. By [25, Chapter 14, Lemma 29] they even meet all inextendible causal smooth curves. According to [8, (9) Remark, (36) Theorem] a spacelike Cauchy surface in Minkowski spacetime is just a causal base, i.e., it is a spacelike [^2] A set is called spacelike according to sec. 1. set which intersects every causal line. It equals the graph of the corresponding $1$-Lipschitz function.
       The basic idea of the proof of the crucial technical result (4) is the application of the Divergence Theorem to a distorted cylindrical surface in $\mathbb{R}^{4}$ (which is $\partial E_{R}$ in the proof of (4)) with a flat bottom base, an achronal top base, and the side constituted by integral curves of the divergence-free $C^{1}$ vector field. As the flux through the side is zero, the fluxes through the bases must be equal. This were the desired result. The crux is that the versions of the Divergence Theorem known to us do not apply to this kind of surface, which obviously is not polyhedral, nor $C^{1}$ as the top base need not be $C^{1}$. Actually, it is not even Lipschitz, but almost Lipschitz. In the sense of definition (3) the set $M$ consists of the interior points of the side and the bases. The complement $M_{0}$ of $M$ is the union of the boundaries of the bases. Here the extension (3) of the Divergence Theorem applies as the exceptional set $M_{0}$ is small enough, i.e. its $3$-dimensional Minkowski content is zero.
(9) Lemma. Let $\tau:\mathbb{R}^{3}\to\mathbb{R}$ and $S:=\{(\tau(x),x):x\in\mathbb{R}^{3}\}$. Then $S$ is a spacelike Cauchy surface if $|\tau(x)-\tau(y)|<|x-y|$ for $x\neq y$ and $\limsup_{|x|\to\infty}|\tau(x)|/|x|<1$ or if a fortiori $\tau$ is $L$-Lipschitz with $L<1$, in which case $S$ is called $L$-achronal.
Proof. Obviously $S$ is spacelike. Assume $(\mathfrak{a}+\mathbb{R}\mathfrak{z})\cap S=\emptyset$ for some $\mathfrak{z}=(1,e)$ with $0<|e|\leq 1$. Let $a_{0}>\tau(a)$. (The case $a_{0}<\tau(a)$ is analogous.) Then by continuity $a_{0}+s>\tau(a+se)$ for all $s$. Hence for $s<-a_{0}$ one has $|\tau(a+se)|/|a+se|>|a_{0}+s|/|a+se|\to 1/|e|\geq 1$ for $s\to-\infty$. This contradicts the assumption $\limsup_{|x|\to\infty}|\tau(x)|/|x|<1$.∎
A $C^{1}$ vector field $\mathfrak{j}$ on $\mathbb{R}^{4}$ is said to satisfy the continuity equation if $\operatorname{div}\mathfrak{j}=0$, i.e., if
$\partial_{0}j_{0}(\mathfrak{x})+\partial_{1}j_{1}(\mathfrak{x})+\,\partial_{2}j_{2}(\mathfrak{x})+\partial_{3}j_{3}(\mathfrak{x})=0$ (4.1)
holds for $\mathfrak{x}\in\mathbb{R}^{4}$.
(10) Theorem. Let the real $C^{1}$ vector field $\mathfrak{j}$ on $\mathbb{R}^{4}$ be bounded and satisfy the continuity equation (4.1). Suppose that $\mathfrak{j}$ is zero or causal future-directed, i.e.
  • $j_{0}(\mathfrak{x})\geq|j(\mathfrak{x})|$ for all $\mathfrak{x}\in\mathbb{R}^{4}$
Then for every spacelike Cauchy surface $S$ being the graph of the corresponding $1$-Lipschitz function $\tau:\mathbb{R}^{3}\to\mathbb{R}$
$\int\big{(}j_{0}(\tau(x),x)-j(\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x=\int j_{0}(0,x)\operatorname{d}^{3}x$ (4.2)
holds. The integrands are nonnegative and the integrals may be infinite, $\nabla\tau$ is measurable a.e. determined.
Proof. By the assumption on $\mathfrak{j}$ and $|\nabla\tau|\leq 1$ the integrands are nonnegative. Let $k:\mathbb{R}^{3}\to\mathbb{R}$ be positive bounded integrable $C^{1}$ like $k(x)=(1+x^{2})^{-2}$.
       (a) The $C^{1}$ vector field $\mathfrak{v}=(v_{0},v):=(j_{0}+k,j)$ is bounded so that its flow is complete, satisfies the continuity equation, and everywhere holds $v_{0}>|v|$. Hence the integral curves $\gamma_{x}$ of $\mathfrak{v}$ determined by $\dot{\gamma}_{x}(s)=\mathfrak{v}(\gamma_{x}(s))$, $\gamma_{x}(0)=(0,x)$, $x\in\mathbb{R}^{3}$ are timelike future directed inextendible. So they meet $S$ just once, i.e., there is exactly one $\sigma(x)\in\mathbb{R}$ such that $\gamma_{x}(\sigma(x))\in S$. Vice versa the integral curve which starts at the point $\mathfrak{y}\in S$ meets $\{0\}\times\mathbb{R}^{3}$. This proves that $\mathfrak{h}:\mathbb{R}^{3}\to\mathbb{R}^{4}$, $\mathfrak{h}(x):=\gamma_{x}(\sigma(x))$ is injective with $\mathfrak{h}(\mathbb{R}^{3})=S$.
       (b) The claim is that $\mathfrak{h}$ is locally Lipschitz. By [13, Proposition 37], $\Phi:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}^{4}$, $\Phi(s,x):=\gamma_{x}(s)$ is a diffeomorphism. Write $\Phi=:(\varphi,\phi)$.
       From the implicit function theorem [20] applied to the function $f:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}$, $f:=\varphi-\tau\circ\phi$, it follows that $\sigma:\mathbb{R}^{3}\to\mathbb{R}$ is continuous and hence that $\mathfrak{h}$ is continuous.
       Indeed, since $f(s,x)=0$ is uniquely solved by $s=\sigma(x)$ for every $x\in\mathbb{R}^{3}$ it remains to verify that $f(\cdot,x)$ is injective for every $x\in\mathbb{R}^{3}$. Let $f(s,x)=f(s^{\prime},x)$. This implies $|\varphi(s,x)-\varphi(s^{\prime},x)|\leq|\phi(s,x)-\phi(s^{\prime},x)|$. Since $\gamma_{x}$ is timelike, $s=s^{\prime}$ follows.
       Now we show that $\mathfrak{h}$ is even locally Lipschitz. Put $\mathfrak{G}:=(\varphi,\operatorname{id}_{\mathbb{R}^{3}})$. As $\partial_{s}\varphi(s,x)=v_{0}(\Phi(s,x))>0$ one infers that $\mathfrak{G}$ is an diffeomorphisms. It follows $g(h_{0}(x),x)=h(x)$ for $g:=\phi\circ\mathfrak{G}^{-1}$.
       Fix $x,y\in\mathbb{R}^{3}$. Define $\beta:[0,1]\to\mathbb{R}^{3}$, $\beta(u):=g(\mathfrak{z}(u))$ with $\mathfrak{z}(u):=\big{(}(h_{0}(y),y)-(h_{0}(x),x)\big{)}u+(h_{0}(x),x)$. Note $\beta(1)=h(y)$, $\beta(0)=h(x)$ and remind $\mathfrak{h}(\mathbb{R}^{3})=S$.
       Then $|h_{0}(y)-h_{0}(x)|=|\tau(h(y))-\tau(h(x))|\leq|h(y)-h(x)|$, whence it suffices to show that $h$ is locally Lipschitz, and further $|h(y)-h(x)|=|\beta(1)-\beta(0)|\leq|\beta^{\prime}(u_{0})|$ for some $u_{0}$ by the mean value theorem.
       Let $D_{\text{time}}g$ denote the $\mathbb{R}^{3}$-valued function obtained by differentiating each component of $g$ with respect to its first variable, and let $D_{\text{space}}g$ denote the $3\times 3$ matrix-valued function given by the Jacobian of $g$ with respect to its last three variables. Then $\beta^{\prime}(u)=D_{time}g(\mathfrak{z}(u))\big{(}(h_{0}(y)-h_{0}(x)\big{)}+D_{space}g(\mathfrak{z}(u))\big{(}y-x\big{)}$, whence
$|h(y)-h(x)|\leq|D_{time}g(\mathfrak{z}(u_{0})|\,|h(y)-h(x)|+||D_{space}g(\mathfrak{z}(u_{0}))||\,|y-x|$
Obviously $||D_{space}g||$ is locally bounded on $\mathbb{R}^{4}$. The time derivative yields $D_{time}g(\mathfrak{z})=(\frac{1}{v_{0}}v)(\mathfrak{G}^{-1}(\mathfrak{z}))$, whence $|D_{time}g(\mathfrak{z})|<1$ for all $\mathfrak{z}\in\mathbb{R}^{4}$. Hence $||D_{space}g||/\big{(}1-|D_{time}g|\big{)}$ is locally bounded on $\mathbb{R}^{4}$; for bounded $D\subset\mathbb{R}^{4}$ let $C_{D}<\infty$ be a bound on $D$.
       Let $K\subset\mathbb{R}^{3}$ be a compact ball. The image of $K$ under the continuous map $(h_{0},\operatorname{id}_{\mathbb{R}^{3}})$ is contained in a compact ball $D\subset\mathbb{R}^{4}$, which is convex. Hence for $x,y\in K$ one has $\mathfrak{z}(u)\in D$ and therefore in conclusion $|h(y)-h(x)|\leq C_{D}|y-x|$.
       (c) Let $0<R<\infty$. Put $A_{R}:=\{x\in\mathbb{R}^{3}:|x|=R\}$ and $B_{R}:=\{x\in\mathbb{R}^{3}:|x|<R\}$. Suppose that $\sigma(x)>0$ for $|x|<R$. Then
$E_{R}:=\bigcup_{x\in B_{R}}\{\gamma_{x}(s):0<s<\sigma(x)\}$
is open in $\mathbb{R}^{4}$ with $\partial E_{R}=M\cup M_{0}$, where $M:=\{0\}\times B_{R}\cup L_{R}\cup\mathfrak{h}(B_{R})$ with $L_{R}:=\bigcup_{x\in A_{R}}\{\gamma_{x}(s):0<s<\sigma(x)\}$, and $M_{0}:=\{0\}\times A_{R}\cup\mathfrak{h}(A_{R})$.
       Now the proof of the basic idea in (d)-(f) follows. One verifies that the divergence theorem (3) applies to $n=4$, $E=E_{R}$ showing $\mathcal{H}^{3}(\partial E_{R})<\infty$ by (d), (e), and $\mathcal{M}^{3}(M_{0})=0$ by (f), (3). It is decisive that the integral over $L_{R}$ is zero as $L_{R}$ is constituted of integral curves.
       (d) Let $C\subset\mathbb{R}^{3}$ be bounded Borel. Then $\mathcal{H}^{3}(\{0\}\times C)<\infty$, $\mathcal{H}^{3}(\operatorname{graph}\tau|_{C})<\infty$ by [21, Theorem 8.1].
       (e) Let $D$ be a bounded Borel subset of the tube $T_{R}:=\{\gamma_{x}(s):s\in\mathbb{R},|x|=R\}$. Then $\mathcal{H}^{n-1}(D)<\infty$. Indeed, $A_{R}$ is covered by $A:=\{x\in A_{R}:|x_{1}|\leq R/3,|x_{2}|\leq R/3\}$ and finitely many rotations of $A$ around the origin. Hence it suffices to show that $\mathcal{H}^{n-1}(D^{\prime})<\infty$ for $D^{\prime}:=\{\gamma_{x}(s):|s|\leq S,x\in A\}$ and $0<S<\infty$.
       Recall the diffeomorphism $\Phi:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}^{4}$. Hence $f:=\mathbb{R}\times]-R/2,R/2[\times]-R/2,R/2[\to\mathbb{R}^{4}$, $f(s,x_{1},x_{2}):=\Phi\big{(}s;x_{1},x_{2},\sqrt{R^{2}-x_{1}^{2}-x_{2}^{2}}\,\big{)}$ is injective, $C^{1}$ whence Lipschitz on compact sets, and $f([-S,S]\times[-R/3,R/3]\times[-R/3,R/3])=D^{\prime}$. The claim holds by [21, Theorem 8.1].
       (f) Let $C\subset\mathbb{R}^{3}$ be a compact Lebesgue null set. Then $\mathcal{H}^{3}(\{0\}\times C)=0$ and $\mathcal{H}^{3}(\mathfrak{h}(C))=0$ hold by [21, Theorem 8.1]. Indeed, the former is obvious, the latter holds true as $\mathfrak{h}$ is Lipschitz on $C$ being compact.
       (g) Now let $\eta\in C_{c}^{1}(\mathbb{R}^{4})$ with $0\leq\eta\leq 1$ and $\eta|_{\overline{E}_{R}}=1$. Put $\phi:=\eta\mathfrak{v}$. Then (3) yields $0=\int_{M}\mathfrak{v}\,\nu_{E_{R}}\operatorname{d}\mathcal{H}^{3}$ as $\mathfrak{v}$ satisfies the continuity equation. Moreover the integration over $L_{R}$ yields 00 as the integrand is 00. So using [21, Theorem 9.1] it follows $\int_{\varpi(\mathfrak{h}(B_{R}))}\big{(}v_{0}(\tau(x),x)-v(\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}x^{3}=\int_{B_{R}}v_{0}(0,x)\operatorname{d}x^{3}$.
       (h) Finally, as explained in [13, Proposition 37], the condition $\sigma(x)>0$ for $|x|<R$ in (c) can be removed and the limit $R\to\infty$ carried out. One ends up with (4.2).∎
Under the assumptions made on the vector field $\mathfrak{j}$, the result in (4) states that the future-directed flux passing through a spacelike Cauchy surface is the same for all these surfaces. Under the additional assumption (4)(b) on $\mathfrak{j}$ the result in (4) is extended to all maximal achronal sets containing the origin simply by flatten them becoming $\gamma$-achronal for $\gamma<1$.
(11) Lemma. Let the real $C^{1}$ vector field $\mathfrak{j}$ be bounded and satisfy the continuity equation (4.1). Suppose that
  • (a)
    $j_{0}(\mathfrak{x})\geq|j(\mathfrak{x})|$ for $\mathfrak{x}\in\mathbb{R}^{4}$
  • (b)
    $j_{0}(\mathfrak{x})\leq C(1+|x|)^{-N}$ for $|x|\geq|x_{0}|$ with some constants $N>3$ and $C<\infty$
Then for every maximal achronal set $\Lambda$ with $0\in\Lambda$
$\int\big{(}j_{0}(\tau(x),x)-j(\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x=\int j_{0}(0,x)\operatorname{d}^{3}x$ (4.3)
holds. Here $\operatorname{graph}\tau=\Lambda$ with $\tau(0)=0$. The integrands are nonnegative and the integrals are finite.
Proof. Due to (a) and $|\nabla\tau|\leq 1$ the integrands are nonnegative. Note $|\tau(x)|=|\tau(x)-\tau(0)|\leq|x-0|=|x|$. Hence, by (b), $\big{|}j_{0}(\tau(x),x)-j(\tau(x),x)\nabla\tau(x)\big{|}\leq 2j_{0}({\tau(x)},x)\leq 2C_{N}(1+|x|)^{-N}$ with $N>3$, whence the integrals are finite.
       Let $10<\gamma<$ 1. Then $\mathfrak{j}(\gamma x_{0},x)\to\mathfrak{j}(x_{0},x)$ for $\gamma\to 1$ by continuity. Note that $\gamma\tau$ is $\gamma$-Lipschitz. Hence by (4) the corresponding maximal achronal set is a spacelike Cauchy surface. Therefore by (4) equation (4.3) holds for $\gamma\tau$ in place of $\tau$. Note that still $|\gamma\tau(x)|\leq|x|$. Thus, by (b), the map $\mathbb{R}^{3}\ni x\mapsto 2C_{N}(1+|x|)^{-N}$ is an integrable majorant uniform with respect to $\gamma$, whence the claim by dominated convergence.∎
5 Covariant achronal localization out of covariant conserved current
Let $\mathcal{H}$ be a separable Hilbert space. Let $\mathcal{B}^{ach}$ denote the family of Borel subsets $\Delta$ of $\mathbb{R}^{4}$, which are achronal.
(12) Definition. Let $T(\Delta)$ for $\Delta\in\mathcal{B}^{ach}$ be a nonnegative bounded operator on $\mathcal{H}$. Suppose $T(\emptyset)=0$ and $\sum_{n}T(\Delta_{n})=I$ for every sequence $(\Delta_{n})$ of mutually disjoint sets in $\mathcal{B}^{ach}$ such that $\bigcup_{n}\Delta_{n}$ is maximal achronal. Then the map $T$ is called an achronal localization (AL).
       Let $W$ be a unitary representation of $\tilde{\mathcal{P}}$. Then the AL $T$ is said to be (Poincaré) covariant by means of $W$ if $T(g\cdot\Delta)=W(g)T(\Delta)W(g)^{-1}$ holds for $g\in\tilde{\mathcal{P}}$ and $\Delta\in\mathcal{B}^{ach}$.
       As mentioned the meaning of $T$ is that $\langle\phi,T(\Delta)\phi\rangle$ is the probability of localization of the quantum mechanical system in the spacetime region $\Delta$ if the system is in the state $\phi$.
       There exist AL with the localization operators $T(\Delta)$ being orthogonal projections [8, (22) Theorem]. In this case the localization operators commute. However, a quantum mechanical system localized by a projection valued AL necessarily does not have a semi-bounded energy operator. This no-go result following from Hegerfeldt’s well-known theorem [17] [^3] See also [18] and D.Buchholz and J.Yngvason’s comment also on Hegerfeld’s achievements [2]. regards a first type of Einstein causality requirement whose modern generalized reformulation [6, 8] we shall present below.
       The notion of localization of the above definition seems appropriate to describe measurement processes where a quantum system is absorbed by the apparatus and no further localization measurments can be performed on it. This is because, in case of subsequent measurements a second type of Einstein’s causality requirement essentially regarding the no-signaling condition has to be considered: The statistics of the measurement results $\Delta$ must be the same, regardless of whether a non-selective measurement was made in $\Delta^{\prime}$ or not.
To comply with this requirement, under some popular assumptions about the post measurement state, $T(\Delta)$ and $T(\Delta^{\prime})$ should commute if $\Delta$ and $\Delta^{\prime}$ cannot be joined by causal curves. [^4] If we explicitly assume that $T(\Delta)$ and $T(\Delta^{\prime})$ are projectors and suppose that the post-measurement state of a non-selective measurement of $T(\Delta^{\prime})$ and $I-T(\Delta^{\prime})$ is obtained with the Lüders-von Neumann projection postulate for every initial state, the no-signaling requirement easily implies that $T(\Delta)$ and $T(\Delta^{\prime})$ commute. If $T(\Delta)$ and $T(\Delta^{\prime})$ are effects, the no-signaling condition is guaranteed if, more weakly, the respective Kraus operator commute. However, if we also assume, as is common in local quantum physics, that all relevant operators (including Kraus ones) are organized in local von Neumann algebras (here associated to the causal completions of $\Delta$ and $\Delta^{\prime}$), the Kraus operator of $T(\Delta)$ must also commute with the adjoint of the Kraus operator of $T(\Delta^{\prime})$. In summary, $T(\Delta)$ and $T(\Delta^{\prime})$ commute. This requirement cannot be fulfilled as a consequence of Malament’s theorem [22] and its modern re-formulations due to Halvorson and Clifton particular [16]. A quick review on these issues related to various facets of Einstein’s causality and localization appears in the introduction of [23].
       The first-type of causality requirement in the modern generalized perspective states that the probability of localization in a region of influence determined by the limiting velocity of light is not less than that in the region of actual localization. The spacetime region $\Delta^{\prime}\subset\mathbb{R}^{4}$ is a region of influence of $\Delta$ if all causal lines, which intersect $\Delta$, meet $\Delta^{\prime}$. Hence the condition imposed by causality on an AL reads
$T(\Delta)\leq T(\Delta^{\prime})\>.$ (CC)
In [13], condition CC was proved true in a special case where the sets $\Delta$ belong to smooth Cauchy surfaces and in [8], the final result has been established that an AL satisfies CC in full. As argued in [8], CC even necessitate achronal localization. In fact spacelike localization is not sufficient since CC induces the localization in achronal hyperplanes. This fact is reported in [8] and studied in detail in [9]. An important result is that the concepts of achronal localization AL and representation of the causal logic RCL are equivalent (cf. (7)) and that just the existence of a maximal spacelike set, which is not maximal achronal because of a missing piece of an achronal not spacelike hyperplane (see $P$ in [8, (8) Example],) is the reason behind the missing orthomodularity of the lattice of the causally complete sets generated by the spacelike relation (cf. [6, sec. 11.3]).
       Hence we consider the study and the explicit construction of the achronal localization of the massive scalar boson to be of utmost relevance.
       It is the very principle of causality which lets one think of the probability of localization as a conserved quantity reigned by an associated density current $\mathfrak{J}$. Indeed, by (5) a covariant AL can be constructed by means of a covariant conserved current assuming that on the Euclidean space its zeroth component describes the density of the probability of localization, namely explicitly $J_{0}(\phi;0,x)\geq 0$, $x\in\mathbb{R}^{3}$, and
$\langle\phi,T(\Delta)\phi\rangle=\int_{\Delta}J_{0}(\phi;0,x)\operatorname{d}^{3}x$ (5.1)
for Borel $\Delta\subset\mathbb{R}^{3}$.
(13) Definition. Let $W$ be a unitary representation of $\tilde{\mathcal{P}}$ and let $\mathcal{D}$ be a $W$-invariant dense subspace of $\mathcal{H}$. Let $\mathfrak{J}=(J_{0},J)$ be a map from $\mathcal{D}\times\mathbb{R}^{4}$ to $\mathbb{R}^{4}$ such that $\mathfrak{J}(\phi,\cdot)$ is a bounded $C^{1}$ vector field. (i) $\mathfrak{J}$ is a conserved current if the latter satisfies the continuity equation. (ii) $\mathfrak{J}$ is a (Poincaré) covariant current if $\mathfrak{J}\big{(}W(g)\phi,\mathfrak{x}\big{)}=A\cdot\mathfrak{J}\big{(}\phi,g^{-1}\cdot\mathfrak{x}\big{)}$, $g=(\mathfrak{a},A)\in\tilde{\mathcal{P}}$ holds.
The introduction of a suitable dense space $\mathcal{D}$ in (5) takes account of the fact that $\mathfrak{J}$ may be given, as in the case of the massive scalar boson, as an integral operator, which is not defined for all wave functions $\phi$. Moreover, $\mathfrak{J}$ is supposed to be real. Actually, by checking the proof of (5), this is a consequence of $J_{0}$ being real and covariance.
(14) Lemma. A covariant current satisfies $J_{0}(\phi,\mathfrak{x})\geq|J(\phi,\mathfrak{x})|$ for $\mathfrak{x}\in\mathbb{R}^{4}$, $\phi\in\mathcal{D}$, if and only if $J_{0}(\phi;0,x)\geq 0$ for $x\in\mathbb{R}^{3}$, $\phi\in\mathcal{D}$.
Proof.
$J_{0}(\phi;0,x)\geq 0$ for $x\in\mathbb{R}^{3}$, $\phi\in\mathcal{D}$Leftrightarrow $\mathfrak{J}(W(g)^{-1}\phi;0,x)\cdot(1,0,0,0)\geq 0$ for $x\in\mathbb{R}^{3}$, $\phi\in\mathcal{D}$, $g=(\mathfrak{a},A)\in\tilde{\mathcal{P}}$Leftrightarrow $\mathfrak{J}(\phi;\mathfrak{a}+A\cdot(0,x))\cdot(A\cdot(1,0,0,0)\geq 0$ for $x\in\mathbb{R}^{3}$, $\phi\in\mathcal{D}$, $(\mathfrak{a},A)\in\tilde{\mathcal{P}}$Leftrightarrow $\mathfrak{J}(\phi;\mathfrak{x})\cdot\mathfrak{e}\geq 0$ for $\mathfrak{x}\in\mathbb{R}^{4}$, $\phi\in\mathcal{D}$, $\mathfrak{e}^{\cdot 2}=1$, $e_{0}>0$, whence the claim.∎
(15) Definition. A map $q:\mathcal{D}\to\mathbb{R}$ is said to be quadratic if (i) $q(\lambda\phi)=|\lambda|^{2}q(\phi)$ for $\lambda\in\mathbb{C}$, $\phi\in\mathcal{D}$ and (ii) $s_{q}(\phi,\phi^{\prime}):=\frac{1}{4}\sum_{\zeta=1,-1,\operatorname{i},-\operatorname{i}}\zeta q(\zeta\phi+\phi^{\prime})$ for $\phi,\phi^{\prime}\in\mathcal{D}$ is Hermitian sesquilinear. Note that $s_{q}(\phi,\phi)=q(\phi)$ holds by (i), whence (ii) is the polarization identity of $s_{q}$.
(16) Lemma. Assume (5.1) for $\phi\in\mathcal{D}$. Then, for every $x\in\mathbb{R}^{3}$, $J_{0}(\cdot\,;0,x)$ is quadratic.
Proof. Put $q_{x}:=J_{0}(\cdot\,;0,x)$. Note $\langle\phi,T(\Delta)\phi\rangle=\int_{\Delta}q_{x}(\phi)\operatorname{d}^{3}x$ and hence $\langle\phi,T(\Delta)\phi^{\prime}\rangle=\int_{\Delta}s_{q_{x}}(\phi,\phi^{\prime})\operatorname{d}^{3}x$, where $\langle\phi,T(\Delta)\phi^{\prime}\rangle$ is a Hermitian sesquilinear form.
       Hence, as to (5) (i), $\int_{\Delta}\big{(}q_{x}(\lambda\phi)-|\lambda|^{2}q_{x}(\phi)\big{)}\operatorname{d}^{3}x=0$ holds for every $\Delta$, whence the integrand is a.e. zero and by its continuity it is zero everywhere.
       Turn to (5) (ii). Show $s_{q_{x}}(\phi_{1}+\phi_{2},\phi^{\prime})-s_{q_{x}}(\phi_{1},\phi^{\prime})-s_{q_{x}}(\phi_{2},\phi^{\prime})=0$. Indeed, one has $\int_{\Delta}\big{(}s_{q_{x}}(\phi_{1}+\phi_{2},\phi^{\prime})-s_{q_{x}}(\phi_{1},\phi^{\prime})-s_{q_{x}}(\phi_{2},\phi^{\prime})\big{)}\operatorname{d}^{3}x=0$ for every $\Delta$, whence the claim. The remaining properties for $s_{q_{x}}$ regarding a Hermitian sesquilinear form hold analogously.∎
(17) Lemma. Let $\mathfrak{J}$ be covariant. Then $J_{0}(\cdot\,,\mathfrak{x})-J(\cdot\,,\mathfrak{x})\,e$ is quadratic on $\mathcal{D}$ for every $\mathfrak{x}\in\mathbb{R}^{4}$ and $e\in\mathbb{R}^{3}$, $|e|\leq 1$ if and only if $J_{0}(\cdot\,;0,x)$ is quadratic on $\mathcal{D}$ for every $x\in\mathbb{R}^{3}$.
Proof. $J_{0}(\phi;0,x)$ is quadratic regarding $\phi$ for every $x\in\mathbb{R}^{3}$Leftrightarrow $J_{0}(W(g)^{-1}\phi;0,x)$ is quadratic regarding $\phi$ for every $x\in\mathbb{R}^{3}$, $g=(\mathfrak{a},A)\in\tilde{\mathcal{P}}$Leftrightarrow $\mathfrak{J}(W(g)^{-1}\phi;0,x)\cdot(1,0,0,0)$ = $\mathfrak{J}(\phi;\mathfrak{a}+A\cdot(0,x))\cdot(A\cdot(1,0,0,0))$ is quadratic regarding $\phi$ for every $x\in\mathbb{R}^{3}$, $(\mathfrak{a},A)\in\tilde{\mathcal{P}}$Leftrightarrow $\mathfrak{I}(\phi;\mathfrak{x})\cdot\mathfrak{e}$ is quadratic regarding $\phi$ for every $\mathfrak{x}\in\mathbb{R}^{4}$, $\mathfrak{e}^{\cdot 2}=1$, $e_{0}>0$, whence the claim.∎
(18) Lemma. Let $\mathfrak{J}$ be conserved and covariant. Let $\phi\in\mathcal{D}$. Suppose (4)(a),(b) for $\mathfrak{J}(\phi,\cdot)$ and suppose $\int J_{0}(\phi;0,x)\operatorname{d}^{3}x=||\phi||^{2}$. Then
$||\phi||^{2}=\int\big{(}J_{0}(\phi;\tau(x),x)-J(\phi;\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x$
holds for every maximal achronal set with corresponding $1$-Lipschitz function $\tau$.
Proof. By covariance of $\mathfrak{J}$ regarding time translations it follows that $\mathfrak{J}(\phi;\tau(x),x)=\mathfrak{J}(W(-\tau(0))\phi;\tau(x)-\tau(0),x)=\mathfrak{J}(\phi^{\prime};\tau^{\prime}(x),x)$ for $\phi^{\prime}:=W(-\tau(0))\phi$, $\tau^{\prime}:=\tau-\tau(0)$. Hence the right side of (4.3) reads $\int J_{0}(\phi^{\prime};0,x)\operatorname{d}^{3}x=||\phi^{\prime}||^{2}=||\phi||^{2}$.∎
The main result follows.
(19) Theorem. Let the real bounded $C^{1}$ current $\mathfrak{J}$ be conserved covariant. Let $\phi\in\mathcal{D}$. Let $J_{0}$ satisfy
  • (a)
    $J_{0}(\phi;0,x)\geq 0$ for $x\in\mathbb{R}^{3}$, $\int J_{0}(\phi;0,x)\operatorname{d}^{3}x=||\phi||^{2}$, $J_{0}(\cdot\,;0,x)$ is quadratic for every $x\in\mathbb{R}^{3}$
  • (b)
    $J_{0}(\phi,\mathfrak{x})\leq C(1+|x|)^{-N}$ for $|x|\geq|x_{0}|$ with some constants $N>3$ and $C<\infty$ depending on $\phi$
Then there is a unique AL $T$ satisfying for every achronal Borel set $\Delta$
$\langle\phi,T(\Delta)\phi\rangle=\int_{\varpi(\Delta)}\big{(}J_{0}(\phi;\tau(x),x)-J(\phi;\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x$ (1)
where $\tau:\varpi(\Delta)\to\mathbb{R}$ with $\operatorname{graph}\tau=\Delta$. $T$ is covariant.
Proof.
Uniqueness is obvious as $\mathcal{D}$ is dense. Let $\phi\in\mathcal{D}$.
       By (a) and (5), $J_{0}(\phi,\mathfrak{x})\geq|J(\phi,\mathfrak{x})|$.
       Let $\Delta_{0}\in\mathcal{B}^{ach}$. There is a maximal achronal set $\Lambda\supset\Delta_{0}$ being the graph of the corresponding $1$-Lipschitz function $\tau:\mathbb{R}^{3}\to\mathbb{R}$. For Borel $\Delta\subset\Lambda$ put
$\pi_{\phi,\Lambda}(\Delta):=\int_{\varpi(\Delta)}\big{(}J_{0}(\phi;\tau(x),x)-J(\phi;\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x$ (2)
As the integrand is nonnegative, $\pi_{\phi,\Lambda}$ is a $\sigma$-additive measure. By (a) and (5), $\pi_{\phi,\Lambda}(\Lambda)=||\phi||^{2}$. Due to (5), (5), $\phi\mapsto\pi_{\phi,\Lambda}(\Delta)$ is the quadratic form of a bounded Hermitian sesquilinear form on $\mathcal{D}$. Therefore by [13, Lemma 48] there is a bounded operator $T(\Delta)$, $0\leq T(\Delta)\leq I$ with $\pi_{\phi,\Lambda}(\Delta)=\langle\phi,T_{\Lambda}(\Delta)\phi\rangle$.
       Let $\phi\in\mathcal{H}$. Approximating $\phi$ by $\phi_{n}\in\mathcal{D}$ finite additivity of $\Delta\mapsto\langle\phi,T_{\Lambda}(\Delta)\phi\rangle$ follows by continuity. Actually it is $\sigma$-additive. Indeed, let $\Delta_{n}\downarrow_{n}\emptyset$ for Borel $\Delta_{n}\subset\Lambda$. Let $\epsilon>0$. Let $\phi^{\prime}\in\mathcal{D}$ with $||\phi-\phi^{\prime}||\leq\epsilon$. An obvious application of the triangle inequality yields $||\langle\phi,T_{\Lambda}(\Delta_{n})\phi\rangle-\langle\phi^{\prime},T_{\Lambda}(\Delta_{n})\phi^{\prime}\rangle||\leq||\phi-\phi^{\prime}||\,||\phi||+||\phi^{\prime}||\,||\phi-\phi^{\prime}||\leq 2\epsilon||\phi||+\epsilon^{2}\leq C\epsilon$ with $C<\infty$ independent of $\Delta_{n}$. As $\langle\phi^{\prime},T_{\Lambda}(\Delta_{n})\phi^{\prime}\rangle\to 0$, one infers $\langle\phi,T_{\Lambda}(\Delta_{n})\phi\rangle\to 0$, whence the claim.
       So $T_{\Lambda}$ is weakly $\sigma$-additive, which by [27, Theorem 4.28] implies the strong $\sigma$-additivity.
       Note that the definition of $T_{\Lambda}(\Delta_{0})$ via (2) does not depend on $\Lambda$. One may omit the index. Thus $T$ is an AL. It remains to show its covariance, which follows immediately by the subsequent proposition (5)(b).∎
(20) Remark. The assumption (5) (a) can be replaced by
  • (a’)
    $\int J_{0}(\phi;0,x)\operatorname{d}^{3}x=||\phi||^{2}$ and $\langle\phi,T(\Delta)\phi\rangle=\int_{\Delta}J_{0}(\phi;0,x)\operatorname{d}^{3}x$ with nonnegative operators $T(\Delta)$ for bounded Borel $\Delta\subset\mathbb{R}^{3}$
since, arguing as in (5), (a’) implies (a).
Regarding the notations see $(2)$ in the proof of (5).
(21) Proposition. Let $g=(\mathfrak{a},A)\in\tilde{\mathcal{P}}$. Then
       (a) the achronal set $g\cdot\Delta$ is given by
$g\cdot\Delta=\{(\tau_{g}(y),y):y\in\varpi(g\cdot\ \Delta)\}$
for $\tau_{g}(y):=(g\cdot(\tau(x),x))_{0}$ with $x:=S^{-1}(y)$, where
$S:\mathbb{R}^{3}\to\mathbb{R}^{3},\quad S(x):=\varpi\big{(}g\cdot(\tau(x),x)\big{)}$
is a bijection.
       (b) $\pi_{W(g)^{-1}\phi,\Lambda}(\Delta)=\pi_{\phi,g\cdot\Lambda}(g\cdot\Delta)$.
Proof.
(a) Obviously $S$ is surjective. Let $S(x)=S(x^{\prime})$. Hence $\varpi\big{(}g\cdot(\tau(x)-\tau(x^{\prime}),x-x^{\prime})\big{)}=0$ with $\big{(}g\cdot(\tau(x)-\tau(x^{\prime}),x-x^{\prime})\big{)}^{\cdot 2}=\big{(}\tau(x)-\tau(x^{\prime}),x-x^{\prime}\big{)}^{\cdot 2}\leq 0$. Therefore also $(g\cdot(\tau(x)-\tau(x^{\prime}),x-x^{\prime}))_{0}=0$, whence $(g\cdot(\tau(x)-\tau(x^{\prime}),x-x^{\prime}))=0$. This means $(\tau(x)-\tau(x^{\prime}),x-x^{\prime})=0$. So $x=x^{\prime}$. In conclusion $S$ is bijective.
       Note $S(\varpi(\Delta))=\varpi(g\cdot\Delta)$. Therefore $\{(\tau_{g}(y),y):y\in\varpi(g\cdot\ \Delta)\}=\{\big{(}(g\cdot(\tau(x),x))_{0},S(x)\big{)}:x\in\varpi(\Delta)\}=\{g\cdot(\tau(x),x):x\in\varpi(\Delta)\}=g\cdot\Delta$.
       (b) By (5) (ii),
$\pi_{W(g)^{-1}\phi,\Lambda}(\Delta)=\int_{\varpi(\Delta)}\mathfrak{J}\big{(}\phi,g\cdot(\tau(x),x)\big{)}\cdot\big{(}A\cdot(1,\nabla\tau(x))\big{)}\operatorname{d}^{3}x$
equals $\int_{\varpi(\Delta)}\mathfrak{J}\big{(}\phi,\tau_{g}(S(x)),S(x)\big{)}\cdot\big{(}A\cdot(1,\nabla\tau(x))\big{)}\operatorname{d}^{3}x$ which using the image of the Lebesgue measure $\lambda$ becomes
$\int_{S(\varpi(\Delta))}\mathfrak{J}\big{(}\phi,\tau_{g}(y),y\big{)}\cdot\big{(}A\cdot(1,\nabla\tau(S^{-1}(y))\big{)}\operatorname{d}S(\lambda)(y)$
Now recall $S(\varpi(\Delta))=\varpi(g\cdot\Delta)$ and note $\operatorname{d}S(\lambda)/\operatorname{d}\lambda=|\det D\,S^{-1}|=|\det D\,S(S^{-1}(\cdot))|^{-1}$.
       It remains to verify
$(1,\nabla\tau_{g}(y))=|\det D\,S(S^{-1}(y))|^{-1}\,A\cdot(1,\nabla\tau(S^{-1}(y)))$ (*)
which is easy in the case $A\in SU(2)$. So it suffices to check the case $g=\operatorname{e}^{\rho\sigma_{3}/2}$, $\rho\in\mathbb{R}$555 Explicitly $\operatorname{e}^{\,\rho\,\sigma_{3}/2}=\operatorname{diag}(\operatorname{e}^{\rho/2},\operatorname{e}^{-\rho/2})$ acts on $\mathbb{R}^{4}$ by $\left(\begin{array}[]{cccc}\cosh(\rho)&0&0&\sinh(\rho)\\ 0&1&0&0\\ 0&0&1&0\\ \sinh(\rho)&0&0&\cosh(\rho)\end{array}\right)$ for $\rho\in\mathbb{R}$ Put $c:=\cosh\rho$, $s:=\sinh\rho$, $z:=\nabla\tau(x)$, $x=S^{-1}(y)$. The rows of $\big{(}D\,S(x)\big{)}^{-1}$ are $(1,0,0)$, $(0,1,0)$, $\frac{1}{c+sz_{3}}(-sz_{1},-sz_{2},1)$. So the right side of (*) equals $\frac{1}{c+sz_{3}}(c+sz_{3},z_{1},z_{2},cz_{3}+s)$. On the left hand side $\nabla\tau_{g}(y))=\big{(}cz_{1}-\frac{(cz_{3}+s)sz_{1}}{c+sz_{3}},\dots,\frac{cz_{3}+s}{c+sz_{3}}\big{)}$. Hence (*) holds thus accomplishing the proof.∎
6 Covariant achronal localizations of the massive scalar boson
The theory developed so far, in particular the construction of a covariant AL (5), (5), is rather general without requiring additional effort. It refers to any relativistic quantum mechanical system, which is completely determined by a unitary representation $W$ of the Poincaré group. Here we turn to the elementary particle being the massive scalar boson, which is the one-particle system of a free Klein-Gordon quantum field uniquely determined by $W=[m,0,+]$, i.e., the mass $m>0$, the spin 00 and the sign of energy $+$. In the momentum representation with $L^{2}(\mathbb{R}^{3})$ being the space of states $W$ reads [^6] Often one uses the antiunitarily equivalent $\operatorname{e}^{-\operatorname{i}\mathfrak{a}\cdot\,\mathfrak{p}}$.
$\big{(}W(\mathfrak{a},A)\phi\big{)}(p)=\sqrt{\epsilon(q)/\epsilon(p)}\,\operatorname{e}^{\operatorname{i}\mathfrak{a}\cdot\,\mathfrak{p}}\,\phi(q)$ (6.1)
with $\epsilon(p):=\sqrt{m^{2}+p^{2}}$, $\mathfrak{p}:=(\epsilon(p),p)$, $\mathfrak{q}=(q_{0},q):=A^{-1}\cdot\mathfrak{p}$.
       One recalls that the localizability of the massive scalar boson in Euclidean space is described by a Euclidean covariant normalized POVM $T$ on the Borel sets of $\mathbb{R}^{3}$, called a POL (Positive Operator Localization) in [7].
       In the momentum representation by [7, (6.1), (11) Theorem] one has
  • $\langle\phi,T(\Delta)\phi\rangle=\int_{\Delta}J_{0}(\phi,x)\operatorname{d}^{3}x$
i.e., (5.1) holds, where the density of the probability of localization $J_{0}$ is given by
  • $J_{0}(\phi,x)=(2\pi)^{-3}\int\int\,\textsc{k}(k,p)\operatorname{e}^{\operatorname{i}(p-k)x}\overline{\phi(k)}\phi(p)\,\operatorname{d}^{3}k\,\operatorname{d}^{3}p$
for $\phi\in C_{c}$, i.e., continuous with compact support. Here k is any measurable normalized (i.e., $\textsc{k}(p,p)=1)$ rotational invariant positive definite separable kernel k on $\mathbb{R}^{3}\setminus\{0\}$.
       The aim is to extend $T$ to an AL which is Poincaré covariant under the representation $W$ (6.1) of the massive scalar boson following the considerations in sec. 5. Petzold and collaborators [15] show that $J_{0}$ is the zero component of a covariant conserved four-vector current $\mathfrak{J}:=(J_{0},J)$ if and only if
$\mathfrak{J}(\phi,\mathfrak{x})=(2\pi)^{-3}\int\int\,\mathfrak{K}(k,p)\operatorname{e}^{\operatorname{i}\big{(}(\epsilon(k)-\epsilon(p))x_{0}-(k-p)x\big{)}}\overline{\phi(k)}\phi(p)\,\operatorname{d}^{3}k\operatorname{d}^{3}p$ (6.2)
with $\phi\in C_{c}$. Here
$\mathfrak{K}(k,p):=\frac{(\epsilon(k)+\epsilon(p),k+p)}{2\sqrt{\epsilon(k)}\sqrt{\epsilon(p)}}g\big{(}\epsilon(k)\epsilon(p)-kp\big{)}$ (6.3)
where $g:[m^{2},\infty[\to\mathbb{R}$ is continuous with $g(m^{2})=1$ such that the zeroth component $K_{0}$ of $\mathfrak{K}$ is a positive definite kernel on $\mathbb{R}^{3}$ (see also [7, (55) Corollary]).
(22) Definition. $\mathfrak{K}$ in (6.3) is called a causal kernel if its zeroth component is positive definite on $\mathbb{R}^{3}$.
For a thorough analysis of the solutions $g$ see [7]. We mention $|g(t)|<g_{3/2}(t)$ if $t\neq m^{2}$, $g\neq g_{3/2}$, where $g_{r}(t):=(2m^{2})^{r}(m^{2}+t^{2})^{-r}$ for $r\geq 3/2$ denotes the basic series of solutions revealed by [15] and [19].
Henceforth we deal with the conserved covariant currents $\mathfrak{J}$ with causal kernel (6.2). For $\mathcal{D}:=C^{\infty}_{c}(\mathbb{R}^{3})$ the assumptions on $\mathfrak{J}$ in (5) are satisfied. $\mathfrak{J}(\phi,\cdot)$ is even smooth. Moreover $J_{0}$ satisfies (5.1) and hence (5)(a) by (5). Regarding the assumption (5)(b) one has
(23) Lemma. Let $\phi\in C^{\infty}_{c}(\mathbb{R}^{3})$. For $g$ in (6.3) assume $g\in C^{4}([m^{2},\infty[)$. Then (5)(b) holds.
Proof. Let $x_{0}\neq 0$. Put
$F:\mathbb{R}^{3}\times\mathbb{R}^{3}\to\mathbb{R},\quad F(k,p):=\frac{\varepsilon(k)x_{0}-kx}{|x|+|x_{0}|}-\frac{\varepsilon(p)x_{0}-px}{|x|+|x_{0}|}$
$F$ is $C^{\infty}$. Put $\varphi(k,p):=(2\pi)^{-3}K_{0}(k,p)\overline{\phi(k)}\phi(p)$. $\varphi$ is $C_{c}^{4}$. One has
$J_{0}(\phi,\mathfrak{x})=\int\int\operatorname{e}^{\operatorname{i}(|x|+|x_{0}|)F(k,p)}\varphi(k,p)\operatorname{d}^{3}k\operatorname{d}^{3}p$ (*)
We proceed according the proof of [26, Theorem 1.8], which concerns the non-stationary phase method. A positive lower bound (**) of $|\nabla F|$ is crucial.
       Let $K:=\operatorname{supp}\phi$. Put $\beta:=\max\{\frac{|p|}{\varepsilon(p)}:p\in K\}$. Clearly $10\leq\beta<$ 1. Then
$\nabla_{k}F(k,p)=(|x|+|x_{0}|)^{-1}\big{(}\frac{x_{0}}{\epsilon(k)}k-x\big{)}$
whence $|\nabla_{k}F(k,p)|\geq(|x|+|x_{0}|)^{-1}\big{(}|x|-|x_{0}|\frac{|k|}{\epsilon(k)}\big{)}\geq\frac{|x|-\beta|x_{0}|}{|x|+|x_{0}|}$ for $k\in K$. Now assume $|x|\geq|x_{0}|$. Then $|\nabla_{k}F(k,p)|\geq\frac{1-\beta}{2}$ and similarly $|\nabla_{p}F(k,p)|\geq\frac{1-\beta}{2}$ for $p\in K$. It follows
$|\nabla F(k,p)|\geq\frac{1-\beta}{\sqrt{2}}>0\;\text{ for }|x|\geq|x_{0}|,\;(k,p)\in K\times K$ (**)
Note also that the derivatives satisfy $|D_{k}^{\alpha}F(k,p)|\leq 1$, $|D_{p}^{\alpha}F(k,p)|\leq 1$ for $|\alpha|=1$, $|D_{k}^{\alpha}F(k,p)|\leq|D^{\alpha}\varepsilon(k)|$, $|D_{p}^{\alpha}F(k,p)|\leq|D^{\alpha}\varepsilon(p)|$ for $|\alpha|\geq 2$. Moreover, $\operatorname{supp}\varphi\subset K\times K$.
       Repeated integration by parts as in the proof of [26, Theorem 1.8] yields
$J_{0}(\phi,\mathfrak{x})=(|x|+|x_{0}|)^{-n}\int_{K^{2}}\operatorname{e}^{\operatorname{i}(|x|+|x_{0}|)F(v)}\Phi^{(n)}(v)\operatorname{d}^{6}v\;\text{ for }|x|\geq|x_{0}|,\,n\leq 4$
using the notation $v:=(k,p)\in\mathbb{R}^{6}$ and $K^{2}=K\times K\subset\mathbb{R}^{6}$. Here the function $\Phi^{(n)}$ is a sum of products of factors $|\nabla F|^{-2}$, $D^{\alpha}F$ for $1\leq|\alpha|\leq 4$, $D^{\alpha}\varphi$ for $|\alpha|\leq 4$. Hence $\Phi^{(n)}$ is bounded on $K^{2}$ independent of $|x|,|x_{0}|$. The result follows.∎
Note that $g$ from the basic series $g_{r}(t)=(2m^{2})^{r}(m^{2}+t^{2})^{-r}$, $r\geq 3/2$ is even $C^{\infty}$, in particular the distinguished $g=g_{3/2}$.
One summarizes (6.2), (6), (5).
(24) Theorem. Let $\mathfrak{J}$ be a covariant conserved current 777recall (5) with causal kernel for the massive scalar boson. Assume $g\in C^{4}([m^{2},\infty[)$. Then there is an AL $T$ satisfying for every achronal Borel set $\Delta$ and $\phi\in C^{\infty}_{c}(\mathbb{R}^{3})$
$\langle\phi,T(\Delta)\phi\rangle=\int_{\varpi(\Delta)}\big{(}J_{0}(\phi;\tau(x),x)-J(\phi;\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x$
where $\tau:\varpi(\Delta)\to\mathbb{R}$ with $\operatorname{graph}\tau=\Delta$. $T$ is covariant.
We turn to the family of localizations of the massive scalar boson obtained out of its stress energy tensor [23]. For a thorough treatment see [13, sec. 6]. The family is indexed by the normalized future-directed timelike fourvectors $\mathfrak{n}$, i.e. $\mathfrak{n}^{\cdot 2}=1$, $n_{0}>0$. The related currents read still in the momentum representation with $\phi\in C_{c}^{\infty}(\mathbb{R}^{3})$
$\mathfrak{J}_{\mathfrak{n}}(\phi,\mathfrak{x})=(2\pi)^{-3}\int\int\,\mathfrak{K}_{\mathfrak{n}}(k,p)\operatorname{e}^{\operatorname{i}\big{(}(\epsilon(k)-\epsilon(p))x_{0}-(k-p)x\big{)}}\overline{\phi(k)}\phi(p)\,\operatorname{d}^{3}k\operatorname{d}^{3}p$ (6.4)
$\mathfrak{K}_{\mathfrak{n}}(k,p):=\frac{\mathfrak{k}\cdot\mathfrak{n}\;\mathfrak{p}+\mathfrak{p}\cdot\mathfrak{n}\;\mathfrak{k}-(m^{2}+\mathfrak{k}\cdot\mathfrak{p})\;\mathfrak{n}}{2\sqrt{\epsilon(k)}\sqrt{\epsilon(p)}}$ (6.5)
with $\mathfrak{p}:=(\epsilon(p),p)$, $\mathfrak{k}:=(\epsilon(k),k)$.
       One easily checks that $\mathfrak{J}_{\mathfrak{n}}(\phi,\cdot)$ is real smooth bounded conserved and that $J_{\mathfrak{n},0}(\cdot\,;0,x)$ is quadratic for every $x\in\mathbb{R}^{3}$. Also one verifies the covariance
$\mathfrak{J}_{\mathfrak{n}}\big{(}W(g)\phi,\mathfrak{x}\big{)}=A\cdot\mathfrak{J}_{A^{-1}\cdot\mathfrak{n}}(\phi,g^{-1}\mathfrak{x})$ (6.6)
for all $\mathfrak{n}$, $g=(\mathfrak{a},A)$, $\mathfrak{x}$. In addition, for every $\mathfrak{n}$, $\phi$, one has $J_{\mathfrak{n},0}(\phi,;0,\cdot)\geq 0$ and $\int J_{\mathfrak{n},0}(\phi,;0,x)\operatorname{d}^{3}x=||\phi||^{2}$ as shown in [13, (64), Theorem 54]. Lemma (5), (5), and (5) hold for every $\mathfrak{J}_{\mathfrak{n}}$ by the same proofs due to (6.6). Finally, (5)(b) holds for every $J_{\mathfrak{n},0}$ by a proof analogous to that of (6). In summary, (5) applies to $\mathfrak{J}_{\mathfrak{n}}$. It follows
(25) Theorem. For every $\mathfrak{n}$ with $\mathfrak{n}^{\cdot 2}=1$, $n_{0}>0$, there is an AL $M^{\mathfrak{n}}$ satisfying for every achronal Borel set $\Delta$ and $\phi\in C^{\infty}_{c}(\mathbb{R}^{3})$
$\langle\phi,M^{\mathfrak{n}}(\Delta)\phi\rangle=\int_{\varpi(\Delta)}\big{(}J_{\mathfrak{n},0}(\phi;\tau(x),x)-J_{\mathfrak{n}}(\phi;\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x$
where $\tau:\varpi(\Delta)\to\mathbb{R}$ with $\operatorname{graph}\tau=\Delta$. One has the covariance $W(g)M^{\mathfrak{n}}(\Delta)W(g)^{-1}=M^{g\cdot\mathfrak{n}}(g\cdot\Delta)$.
7 Covariant representation of the causal logic for the massive scalar boson
The causal logic $\mathcal{C}$ is the lattice of Borel subsets of $\mathbb{R}^{4}$ which is partially ordered by set inclusion $\subset$ and which is generated and orthocomplemented by achronal separateness, i.e., the relation
$\mathfrak{x}\perp\mathfrak{y}\quad\Leftrightarrow\quad\mathfrak{x}\neq\mathfrak{y}\textnormal{ and }(\mathfrak{x}-\mathfrak{y})^{\cdot 2}\leq 0\>.$ (7.1)
The elements of $\mathcal{C}$ are called causally complete sets. In fact, they satisfy the requirement Δ=($\Delta=(\Delta^{\perp})^{\perp}$ where $M^{\perp}:=\{\mathfrak{x}\in\mathbb{R}^{4}\>:\>\mathfrak{x}\perp\mathfrak{y}\>,\forall\mathfrak{y}\in M\}$ is the causal complement of $M\subset\mathbb{R}^{4}$, and $(M^{\perp})^{\perp}$ is the causal completion of $M$.
       The causal logic is thoroughly studied by Cegla, Jadczyk, in [10] and has been studied further on. $\mathcal{C}$ enjoys some appealing physical properties similar to the ones of the lattice of abstract elementary propositions of a quantum system [10] and see also the more recent works [4, 5]: it is possible to prove that the lattice $\mathcal{C}$ is $\sigma$-complete, irreducible, orthomodular, atomic, atomistic exactly as for a quantum lattice, but fails to satisfy the covering law and separability.
       Since the work of [10] there persists the outstanding question how to construct covariant representations of the causal logic.
(26) Definition. Let $F(M)$ for $M\in\mathcal{C}$ be a bounded nonnegative operator on $\mathcal{H}$. Suppose $F(\emptyset)=0$, and $\sum_{n}F(M_{n})=I$ for every sequence $(M_{n})$ of mutually orthogonal sets in $\mathcal{C}$ such that $\bigvee_{n}M_{n}=\mathbb{R}^{4}$. Then the map $F$ is called a representation of the causal logic (RCL).
       Let $W$ be a unitary representation of $\tilde{\mathcal{P}}$. Then the RCL $F$ is said to be (Poincaré) covariant by means of $W$ if $F(g\cdot M)=W(g)F(M)W(g)^{-1}$ holds for $g\in\tilde{\mathcal{P}}$ and $M\in\mathcal{C}$.
The convergence of the above sum occurs in the weak operator topology (equivalently in the strong operator topology).
By the way a RCL $F$ is a map from the lattice $\mathcal{C}$ to the generalized $\sigma$ effect algebra of effects $E(\mathcal{H})$ on $\mathcal{H}$ [12]. This latter enjoys a weakened form of $\sigma$-complete orthocoplemented lattice structure. $F$ is a homomorphism of this weakened type of structure. In particular, it is easy to prove that a RCL is $\sigma$-additive, order-preserving and orthocomplement-preserving.
Only recently covariant RCL have been constructed. They concern quantum mechanical systems with definite spin $j\in\mathbb{N}_{0}/2$ and mass spectrum $\subset$ ]0,[]0$\infty$[] 0 , ∞ [ of positive Lebesgue measure [8]. Moreover, due to the results in [6], there is evidence that there are also RCL regarding the Dirac particles and the four Weyl fermions. These results have been obtained by group theoretical methods.
       A different approach is to relate RCL to Poincaré covariant conserved (operator) density currents. See the works cited in [3] and [11]. A first concrete step in realizing a representation is done in [11] essentially showing (4.3) for a smooth conserved current with compact support.
       Recently an other idea of localization of the massive scalar boson in causally complete regions is pursued [24]. It uses the method of modular localization from AQFT. Given a Cauchy surface, via the modular localization map, to every state and to the causal completion of every of its Borel subsets a probability of localization is attributed. This is asymptotically additive when increasing the distance between the regions of localization.
       There is the closed relationship (7) between AL and RCL, which is easy to expound. The set of determinacy of $M\subset\mathbb{R}^{4}$ is defined as
$M^{\sim}:=\{\mathfrak{x}:\forall\;\mathfrak{z}\textnormal{ with }\mathfrak{z}^{\cdot 2}>0\;\exists\,s\in\mathbb{R}\textnormal{ with }\mathfrak{x}+s\mathfrak{z}\in M\}\>.$
It consists of all points $\mathfrak{x}$ such that every timelike line through $\mathfrak{x}$ meets $M$. There is the remarkable result that, if $\Delta\in\mathcal{B}^{ach}$, then $\Delta^{\sim}=(\Delta^{\perp})^{\perp}$, whence $\Delta^{\sim}\in\mathcal{C}$ and conversely, if $M\in\mathcal{C}$ then $M=\Delta^{\sim}$ for $\Delta\in\mathcal{B}^{ach}$ maximal achronal in $M$. For details see [10], [6], [8]. This feature of spacetime is the reason for the following one-to-one correspondence.
(27) Proposition. See [8, (19), (20)].
       (a) Let $F$ be an RCL. Set $T(\Delta):=F(\Delta^{\sim})$ for $\Delta\in\mathcal{B}^{ach}$. Then $T$ is an AL. If $F$ is covariant by means of $W$, then so is $T$.
       (b) Let $T$ be an AL. Then there is a unique RCL $F$ with $F(\Delta^{\sim})=T(\Delta)$ for $\Delta\in\mathcal{B}^{ach}$. If $T$ is covariant by means of $W$, then so is $F$.
It suffices to join up (6), (7).
(28) Theorem. To every causal kernel $\mathfrak{K}$ with $g\in C^{4}([m^{2},\infty[)$ there is a unique covariant888recall (7) RCL $F$ for the massive scalar boson such that
$\langle\phi,F(\Delta^{\sim})\phi\rangle=\langle\phi,T(\Delta)\phi\rangle$
holds for $\phi\in C^{\infty}_{c}(\mathbb{R}^{3})$ and every achronal Borel set $\Delta$. Here $\langle\phi,T(\Delta)\phi\rangle$ is given in (6) with $\mathfrak{J}$ from (6.2).
Thus, apparently for the first time, a covariant RCL for a quantum mechanical system with definite mass is achieved.
       Similarly one obtains from (6) a covariant family of RCL related to the stress energy tensor of the massive scalar boson.
(29) Theorem. For every $\mathfrak{n}$ with $\mathfrak{n}^{\cdot 2}=1$, $n_{0}>0$, there is a unique RCL $F^{\mathfrak{n}}$ for the massive scalar boson such that
$\langle\phi,F^{\mathfrak{n}}(\Delta^{\sim})\phi\rangle=\langle\phi,M^{\mathfrak{n}}(\Delta)\phi\rangle$
holds for $\phi\in C^{\infty}_{c}(\mathbb{R}^{3})$ and every achronal Borel set $\Delta$. Here $\langle\phi,M^{\mathfrak{n}}(\Delta)\phi\rangle$ is given in (6). One has the covariance $W(g)F^{\mathfrak{n}}(\Delta)W(g)^{-1}=F^{g\cdot\mathfrak{n}}(g\cdot\Delta)$.
8 Final comments and QFT outlook
All achievements presented in this work concern the notion of localization for a boson massive particle in Minkowski spacetime. We constructed very general notions of achronal localization in complete agreement with requirements of Poincaré covariance and causality. It was thanks to suitable probability conserved currents and an appropriate use of the divergence theorem. Referring to the discussion in sec.5, the first type of causal obstructions against any localization notion, historically arising from Hegerfeldt’s analysis and considered here into a more modern perspective started in [6], were found to be definitely harmless. A fully-fledged notion of achronal, causal, covariant localization is proved to be possible, provided one adopts the modern point of view where, in the general case, observables are described by (families of) POVMs instead of (families of) projection-valued measures, namely selfadjoint operators. No sharply localized probability distributions are permitted in this approach. (Though sharply distributions can be approximated with arbitrary precision [6, 23].) It is interesting noting that, classic notions as the Newton-Wigner position operator, apparently ruled out by the Hegerfeldt theorem, still keep to play a role in our framework as the first-moment of the introduced families of POVMs as established in [13] in a wide generality. This fact gives rise to the standard notion of localization in the rest space of an observer under suitable non-relativistic conditions (large mass limit).
Two types of probability currents, associated to respective notions of localization, have been considered in this work for real massive bosons (see sec.6). A type, extensively studied in [7] within a general framework, was based on the theory of causal kernels. Another type discussed in [23, 13], was constructed in terms of the stress-energy tensor.
The new perspective developed in this paper on the one hand refers to a completely general notion of rest space [7] where localization takes place. This notion is even more general than the one studied in [13], and is represented by (maximal) achronal sets, using the weakest possible regularity condition explicitly required by physics. On the other hand, this generality permitted us to construct, apparently for the first time, an effect-space covariant representation of the lattice of causally complete Borel sets of the spacetime, the logic of spacetime, for a boson particle with definite and strictly-positive value of the mass. A similar type of representation for particles with unsharp mass spectrum was already introduced in [7]. In this view, a particle localized in an (roughly speaking $3D$) achronal region $\Delta$ is the same as a particle localized in its (roughly speaking $4D$) causal completion $\Delta^{\sim}$. Or also in every other achronal region $\Delta^{\prime}$ whose causal completion is the same $\Delta^{\sim}$.
From the perspective of QFT, our approach concerns the one-particle structure of the Fock representation built upon the Poincaré invariant Gaussian vacuum state in Minkowski spacetime, for a massive real Klein-Gordon field. As a consequence, at this stage of the study, it is not directly possible to implement a fully-fledged analysis of local causality in the Haag-Kastler perspective. An evident issue concerns the fact that the projector onto the one-particles Hilbert space does not belong to a local algebra of observables. To overcome this obstruction it is necessary to see all introduced notions as restrictions to the one-particle space of more general notions defined in terms of local algebras of the Klein-Gordon field. The restrictions should be considered as imposed by the choice of the used one-particle states.
Referring to the two types of localizations analyzed in sec.6, this is surely possible concerning the notion of localization constructed out the stress energy tensor [13]. This is because the used stress-energy tensor operator is nothing but the one of the full quantum field averaged against one-particle states. The notion arising from causal kernels [7] needs a deeper analysis. However, it seems clear that, at least in the case of a charged field, a relation with the standard conserved current of the whole quantum field must exist.
Yet, the QFT perspective is the right place where tackling the second type of causal obstructions (see sec.5) against any notion of localization in addition to Hegerfeldt’s theorems. We are referring to the already mentioned Malament [22] and Halvorson-Clifton [16] theorems. These no-go results show that localization operators (either orthogonal projectors or effects) associated to causally separated regions $\Delta$ and $\Delta^{\prime}$ cannot commute, under physically plausible assumptions, as additivity and a bound from below of the generator of time translations. The commutativity requirement as mathematical description of causal independence is an indispensable condition in the local algebras formulation.
All these issues will be tackled elsewhere.
Acknowledgements
V.M. and C. De R. are grateful to Silvano Delladio and Andrea Marchese for useful suggestions and discussions, and acknowledge that this work has been written within the activities of INdAM-GNFM.